### v2

# Optimizing Learning for Robust Hyperbolic Deep Learning in Computer Vision

Abstract
Hyperbolic deep learning has become a growing research direction in computer vision for the unique properties afforded by the alternate embedding space. The negative curvature and exponentially growing distance metric provide a natural framework for capturing hierarchical relationships between datapoints and allowing for finer separability between their embeddings. However, these methods are still computationally expensive and prone to instability, especially when attempting to learn the negative curvature that best suits the task and the data. Current Riemannian optimizers do not account for changes in the manifold which greatly harms performance and forces lower learning rates to minimize projection errors. Our paper focuses on curvature learning by introducing an improved schema for popular learning algorithms and providing a novel normalization approach to constrain embeddings within the variable representative radius of the manifold. Additionally, we introduce a novel formulation for Riemannian AdamW, and alternative hybrid encoder techniques and foundational formulations for current convolutional hyperbolic operations, greatly reducing the computational penalty of the hyperbolic embedding space. Our approach demonstrates consistent performance improvements across both direct classification and hierarchical metric learning tasks while allowing for larger hyperbolic models.
1 Introduction
With the recent rise in the use of hyperbolic manifolds for deep representation learning, there is a growing need for efficient, flexible components that can fully exploit these spaces without sacrificing stability. This has led researchers to focus on two main derivations of hyperbolic space: the Poincaré manifold and the hyperboloid. The Poincaré ball, equipped with a gyrovector space, supports various well-defined operations, including generalized vector addition and multiplication, but it suffers from significant stability issues. On the other hand, the hyperboloid, or Lorentz space, lacks these operations but offers much better operation stability, as demonstrated in the study by Mishne et al. (2022).
To address this gap, previous works have sought to provide Lorentzian definitions for common deep learning operations such as the feed-forward layer (Chen et al., 2022; Dai et al., 2021; Ganea et al., 2018), convolutional layer (Chen et al., 2022; Qu & Zou, 2023; Dai et al., 2021), and MLRs (Bdeir et al., 2024). This increased focus on hyperbolic modeling has led to its gradual integration into computer vision architectures, as detailed in the survey by Mettes et al. (2023). Specifically, the hyperboloid model has been employed as a sampling space for VAEs (Nagano et al., 2019), a decoder space for vision tasks in hybrid settings (Guo et al., 2022; Liu et al., 2020; Khrulkov et al., 2020; Qu & Zou, 2022), and ultimately for fully hyperbolic Lorentzian vision encoders (Bdeir et al., 2024) simultaneously with its Poincaré counterpart (van Spengler et al., 2023).
This paper furthers the development of hyperbolic learning for vision tasks, specifically for the Lorentz manifold. Our primary focus is on the challenge of learning the manifold’s negative curvature. The driving principle behind this, is that the model embeddings may exhibit varying degrees of hyperbolicity depending on the innate hierarchies in the datapoints themselves, the problem task that is being considered, and the specific locations of hyperbolic operation integrations. To accomodate for this, we can adjust the embedding space’s hyperbolic metric to be less or more Euclidean which accounts for the modeling requirements. We also build on the idea of separate manifolds for separate main blocks in the architecture further increasing representative flexibility.
We also recognize that despite recent advances, Lorentz models continue to struggle with issues of high computational costs. We attempt to isolate and alleviate the main factors leading to numerical inaccuracies and computational overhead overall, and more particularly when modeling data in higher-dimensional embedding spaces and when learning the curvatures. Our contributions can then be summed up as:
  1. 1.
    We propose a formulation for Riemannian AdamW and an alternative schema for Riemannian optimizers that accounts for manifold curvature learning.
  2. 2.
    We propose the use of our maximum distance rescaling function to restrain hyperbolic vectors within the representative radius of accuracy afforded by the number precision, even allowing for fp16 precision.
  3. 3.
    We provide a more efficient convolutional layer approach that is able to leverage the highly optimized existing implementations.
  4. 4.
    We empirically show the effectiveness of combining these approaches using classical image classification tasks and hierarchical metric learning problems.
2 Related Work
Hyperbolic Embeddings in Computer Vision
With the success of employing hyperbolic manifolds in NLP models (Zhu et al., 2021; Dhingra et al., 2018; Tifrea et al., 2018) hyperbolic embeddings have extended to the computer vision domain. Initially, many of the works relied on a hybrid architecture, utilizing Euclidean encoders and hyperbolic decoders (Mettes et al., 2023). This was mainly due to the high computational cost of hyperbolic operations in the encoder, as well as the lack of well-defined alternatives for Euclidean operations. However, this trend has begun to shift towards the utilization of fully hyperbolic encoders as can be seen in the hyperbolic Resnets by Bdeir et al. (2024) and van Spengler et al. (2023). Both works offer hyperbolic definitions for 2D convolutional layer, batch normalization layer, and an MLR for the final classification head. Bdeir et al. (2024) even attempt to hybridize the encoder by employing the Lorentz manifold in blocks that exhibit higher output hyperbolicity. While this has led to notable performance improvements, both models suffer from upscaling issues. Attempting to replicate these approaches for larger datasets or bigger architectures becomes much less feasible in terms of time and memory requirements. Instead, our approach places higher focus on efficient components to leverage the beneficial hyperbolic properties of the model while minimizing the memory and computational footprint.
Curvature Learning
Previous work in hyperbolic spaces has explored various approaches to curvature learning. In their studies, Gu et al. (2018) and Giovanni et al. (2022) achieve this by using a radial parametrization that implicitly models variable curvature embeddings under an explicitly defined, fixed 1-curve manifold. This method enables them to simulate K-curve hyperbolic and spherical operations under constant curvature for the mixed-curve manifold specifically, a combination of the Euclidean, spherical, and Poincaré manifold. Other approaches, such as the one by Kochurov et al. (2020), simply set the curvature to a learnable parameter but do not account for the manifold changes in the Riemannian optimizers. This leads to hyperbolic vectors being updated with mismatched curvatures and others being inaccurately reprojected, resulting in instability and accuracy degradation. Additionally, some methods, like the one by Kim et al. (2023), store all manifold parameters as Euclidean vectors and project them before use. While this approach partially mitigates the issue of mismatched curvature operations, it remains less accurate and more computationally expensive. In comparison, our proposed optimization schema maintains the parameters on the manifold and optimizes them directly by performing the necessary operations to transition between the variable curvature spaces.
Metric Learning
Metric learning relies on the concept of structuring the distribution in the embedding space so that related data points are positioned closer together, while less related points are placed further apart. To facilitate this process, numerous studies have introduced additional loss functions that explicitly encourage this behavior. Contrastive losses, for instance, operate on pairs of data points and propose a penalty that is proportional to the distances between negative pairs and inversely proportional to the distance between positive pairs (Chopra et al., 2005). Triplet loss extends this idea by considering sets of three points: an anchor, a positive sample, and a negative sample (Wang et al., 2014). Instead of changing the distances between points absolutely, it ensures that the distance between the anchor and the positive sample is less than the distance between the anchor and the negative sample, plus a margin, thus enforcing a relational criterion.
These approaches have also been adapted to hierarchical problem settings under hyperbolic manifolds (Yang et al., 2022; Kim et al., 2023). Notably, Kim et al. (2023) developed a method for learning continuous hierarchical representations using a deep learning, data-mining-like approach that relies on the innate relationships of the embeddings rather than their labels. They employ a proxy-based method that models the data on the Poincaré ball, facilitating a more natural extension to hierarchical tasks. Building on this, we extend the approach by modeling the loss function in the Lorentz manifold and incorporating a learnable curvature to better handle data with varying levels of hierarchy.
3 Methodology
Hyperbolic embeddings offer significant advantages for computer vision tasks that benefit from hierarchical structure modeling, but their effectiveness is hindered by practical challenges, including mathematical inconsistencies in curvature learning, the absence of a Riemannian AdamW optimizer, stability issues with float32 precision, and high computational costs from inefficient convolutions.
To address these issues, the following sections introduce four key advancements: a more stable curvature learning method, a Riemannian AdamW optimizer, a normalization scheme for improved float32 precision stability, and efficient hyperbolic convolutions to reduce memory and computational costs.
3.1 Background
The hyperbolic space is a Riemannian manifold with a constant negative sectional curvature $c<0$. There are many conformal models of hyperbolic space but we focus our work on the hyperboloid, or Lorentz manifold. The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}\in\mathbb{R}^{n+1}~{}|~{}\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=\frac{-1}{K},~{}x_{t}>0\}$ where $\frac{-1}{K}=c$, and with the Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\mathrm{diag}(-1,1,\dots,1)$. This models the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}=[\sqrt{K},0,\cdots,0]^{T}$. We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector, the space dimension $\bm{x_{s}}$. The Lorentzian inner product then becomes $\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}=\bm{x}^{T}\mathrm{diag}(-1,1,\cdots,1)\bm{y}.$ We now define the common hyperbolic operations in the Lorentz space.
Distance
Distance in hyperbolic space is the magnitude of the geodesic forming the shortest path between two points. Let $\bm{x},\bm{y}\in\mathbb{L}^{n}_{K}$, the distance between them is given by $d_{\mathbb{L}}(\bm{x},\bm{y})=\sqrt{K}\cosh^{-1}(\frac{-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}{K})$. We also define the square distance by Law et al. (2019) as $d^{2}_{\mathbb{L}}(\bm{x},\bm{y})=||\bm{x}-\bm{y}||_{\mathbb{L}}^{2}=-2K-2\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$.
Exponential and Logarithmic Maps
Seeing as the Lorentz space is a Riemannian manifold, it is locally Euclidean. This can best be described through the tangent space $\mathbb{T}_{\bm{x}}\mathcal{M}$, a first-order approximation of the manifold at a given point $\bm{x}$. The exponential map, $\exp^{K}_{\bm{x}}\left(\bm{z}\right):\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}\rightarrow\mathbb{L}^{n}_{K}$ is then the operation that maps a tangent vector $\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}$ onto the manifold through $\exp_{\bm{x}}^{K}(\bm{z})=\cosh(\alpha)\bm{x}+\sinh(\alpha)\frac{\bm{z}}{\alpha},\text{ with }\alpha=\sqrt{1/K}||\bm{z}||_{\mathbb{L}},~{}~{}||\bm{z}||_{\mathbb{L}}=\sqrt{{\langle\bm{z},\bm{z}\rangle}_{\mathbb{L}}}.$ The logarithmic map is the inverse of this mapping and can be described as $\log_{\bm{x}}^{K}(\bm{y})=\frac{\cosh^{-1}(\beta)}{\sqrt{\beta^{2}-1}}\cdot(\bm{y}-\beta\bm{x}),\text{ with }\beta=-\frac{1}{K}\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}.$
3.2 Riemannian Optimization
On the Stability of Curvature Learning
In the hyperbolic learning library GeoOpt, Kochurov et al. (2020) introduce the curvature of hyperbolic space as a learnable parameter. However, no subsequent work has fully leveraged this feature, and our empirical testing reveals that this approach often leads to instability and performance decline. We attribute these issues to the simplistic implementation of the curvature updates, which neglects to update the hyperbolic operations after updating the curvature.
In the current Riemannian optimization approach, Euclidean gradients and their momentums are projected onto the Riemannian manifold and then used to update the hyperbolic parameters. These operations inherently depend on the curvature of the space. However, when the curvature is adjusted during learning, the projections, gradients, and hyperbolic parameters can become inconsistent, as they mix the projection and update operations from the new curvature with vectors defined under the previous curvature.
To illustrate this, Figure 1 shows two manifolds with different curvatures. It is evident that points defined on the first manifold do not lie on the second, and the tangent planes of the first manifold are not aligned with those of the second. Thus, using the new curvature operations on points from the old model is not mathematically sound practice. This inconsistency introduces instability, undermining the reliability of curvature learning and often leading to invalid values during the parameter updates. To address this, we propose a method that tracks both the old and new curvature values throughout updates. Specifically, all updates are applied under the old curvature, after which the hyperbolic vectors are mapped from the old manifold to the new one.
While GeoOpt does not address the optimization challenges involved in learning manifold curvature, it does implement an ”N-stabilize step”. This step aims to reduce instability by recalculating the time component of the hyperbolic parameters every N-steps to ensure they remain on the manifold. However, this approach fails to prevent invalid parameter values before the update steps and suffers from high inaccuracies due to the mathematically flawed parameter optimization.
Moreover, the time re-calculation projection used in the stabilize step alters the relative magnitudes of the hyperbolic vectors and gradients, and even the directions of the momentums, which can degrade performance and lead to training instability. Instead, we recommend projecting points onto the tangent space at the old origin using the logarithmic map and then projecting back after the curvature update. Gradients and their momentum can be similarly projected or parallel transported to the tangent space. Parallel transport is chosen here because it preserves the vectors’ directions which is important for the optimization procedure.
Additionally, when the curvature changes, the tangent space at the origin remains relatively stable, as shown in Figure 1. Moving onto the new manifold is then a simple exponential map and parallel transport back operation. This preserves the vector norms and their hyperbolic distances to the origin since $D(\exp^{K_{1}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{1}}=||y||=D(\exp^{K_{2}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{2}}$ where $\bm{y}\in\mathcal{T}_{\overline{\bm{0}}}\mathcal{M}$. No changes are needed for the tangent vector during the move between tangent spaces of different curvatures since these spaces are parallel, making the transition a simple translation along the time axis.
(a) At a random point
(b) At the origin
Figure 1: Tangent planes of a hyperboloid with curvature -1 relative to another hyperboloid with curvature -0.7. Tangential properties between manifolds are better respected at the origin where tangents remain parallel.
It is important to emphasize that our proposed optimization scheme is compatible with existing curvature learning and meta-learning methods. Rather than being an alternative, it serves as a guideline for updating manifold parameters during curvature changes, aimed at maintaining learning stability throughout the process.
Riemannian AdamW Optimizer
Recent works, particularly with transformers, commonly use the AdamW optimizer introduced by Loshchilov & Hutter (2019). As of current, there is no established Riemannian variant of this optimizer. We attempt to derive AdamW for the Lorentz manifold and argue a similar approach could be extended to the Poincaré ball. The main difference between AdamW and Adam is the direct weight regularization. This is more difficult to perform in the Lorentz space given the lack of an intuitive subtraction operation on the manifold. To resolve this, we reframe the parameter regularization in the original AdamW as a weighted centroid with the origin $O$
$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda O$
where $\gamma$ is the learning rate and $\lambda$ is the weight decay value. The centroid operation is well defined in the Lorentz space which now allows for a direct translation of AdamW. The regularization schema becomes:
$\bm{\theta_{t}}=\begin{dcases}\mu_{\mathbb{L}}([\bm{\theta_{t-1}},\overline{\bm{0}}],\bm{\nu}=[1-\gamma\lambda,\gamma\lambda])&\quad\textrm{if }\bm{\theta}\in\mathbb{L}\\ \bm{\theta_{t-1}}-\gamma\bm{\theta_{t-1}}\lambda&\quad\textrm{otherwise}\end{dcases}$
where $\mu_{\mathbb{L}}$ is the Lorentz centroid, and $\bm{\nu}$ are the weights. By removing the later gradient decay and introducing this operation, we adapt AdamW for use in the Lorentz space.
Algorithm 1 Curvature Learning Aware Optimization Given parameters $\bm{\theta}=[\bm{\theta}_{euclid},\bm{\theta}_{\mathbb{L}},\bm{\theta}_{K}]$ and gradients denoted $\mathcal{G}$ procedure Optimizer Step($\bm{\theta}$)      for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do          Vanilla Riemannian Optimizer Step      end for      for $\bm{p}$ in $\bm{\theta}_{euclid}\cup\bm{\theta}_{K}$ do          Vanilla Eucldiean Optimizer Step      end for      MoveParameters() end procedure procedure Move Parameters      for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do          $\bm{\mathcal{G}}_{temp}=\mathrm{PT}^{K_{t-1}}_{{\bm{p}}\rightarrow{\bm{\overline{\bm{0}}}_{t-1}}}\left({\bm{\mathcal{G}}}\right)$          $\bm{z}=\log_{\bm{\overline{\bm{0}}}_{t-1}}^{K_{t-1}}(\bm{p})$          $\bm{p}=\exp_{\overline{\bm{0}}_{t}}^{K_{t}}(\bm{z})$          $\bm{\mathcal{G}}=\mathrm{PT}^{K_{t}}_{{\bm{\overline{\bm{0}}}_{t}}\rightarrow{\bm{p}}}\left({\bm{\mathcal{G}}_{temp}}\right)$      end for end procedure
Maximum Distance Rescaling
[…]
$D^{K}_{\overline{\bm{0}}_{max}}=arccosh(\frac{x_{t_{max}}}{\sqrt{K}})\cdot\sqrt{K}$ (1)
Under Float32 precision, and to account for values of $K<1$ we use a limit value of $x_{t_{max}}=2\cdot 10^{3}$. When projected onto the tangent space of the origin, and with $K=1$, this translates to $||\log^{1}_{\overline{\bm{0}}}(\bm{x})||=D^{1}_{\overline{\bm{0}}_{max}}=9.1$. This value changes considerably as the value of $K$ changes. Vectors outside this radius lead to instability and performance degradation due to inaccurate approximation. This problem is only exacerbated as the dimensionality of the hyperbolic vector increases. Higher dimensional vectors tend to have larger norms which limits hyperbolic models’ abilities to scale up.
To constrain hyperbolic vectors within a specified maximum distance, either a normalization function or a parameter clipping method is required. Parameter clipping can be challenging to train, as it may lead to information loss and introduce non-smooth gradients. On the other hand, common normalization functions like tanh and the sigmoid function tend to saturate quickly, limiting their effectiveness as seen in the sigmoid implementation by Chen et al. (2022). To address these issues, we introduce a modified scaling function, designed to provide finer control over both the maximum values and the slope of the curve. A visualization of this function is provided in Figure 3, and the formulation is presented below:
$\bm{y}_{rescaled}=\frac{\bm{y}}{\lVert\bm{y}\rVert}\cdot m\cdot tanh(\lVert\bm{y}\rVert\cdot\frac{atanh(0.99)}{s\cdot m})$ (2)
where $y\in\mathbb{R}^{d}$, $m$ is our desired maximum value, and $s$ controls the slope of the curve. We now have a maximum distance value to adhere to and a flexible distance normalizing function. To apply this to the hyperbolic vectors, we suggest performing the scaling on the tangent plane of the origin. However, this is an expensive operation to perform often, as such we derive in appendix A.2 the equivalent factorized form for the scaling of the space values:
${\bm{x}_{s_{rescaled}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}}$ (3)
where $D(\bm{x},\overline{\bm{0}})_{rescaled}^{k}$ is the scaled distances by plugging in $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Eq.2. The time component of the vector is then recalculated based on the norm of its spatial components. This gives us a complete scaling operation for the Lorentz space. We apply this tanh scaling when moving parameters across different manifolds. This includes transitions from the Euclidean space to the Lorentz space, as well as between Lorentz spaces of different curvatures. Additionally, we use this scaling after Lorentz Boosts and direct Lorentz concatenations (Qu & Zou, 2022). We also incorporate it following the variance-based rescaling in the batch normalization layer, since variance adjustments during this operation can push points outside the radius during the operation and often lead to invalid values.
3.3 Towards Efficient Architectural Components
Lorentz Convolutional Layer
In their work, Bdeir et al. (2024) proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al. (2022). This approach ensured that the convolution outputs remained on the hyperboloid. However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions.
To address this issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al. (2021), which decomposes the transformation into a Lorentz boost and a Lorentz rotation. Using this definition, we replace the matrix multiplication employed by Bdeir et al. (2024) for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost. Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
A rotation is any operation that modifies the directions of transformed vectors while preserving their magnitudes. This can be achieved using a matrix $\bm{\hat{W}}\in\mathbb{S}$. Where
$\mathbb{S}(n^{\prime},n)={\bm{M}\in\mathbb{R}^{(n^{\prime}\cdot n)}:\bm{M}^{T}\bm{M}=\bm{I}}$ (4)
represents the Stiefel manifold. The condition $n\leq n^{\prime}$ ensures that the columns of the weight matrix are orthogonal, thereby preserving the norms along the rows of the input matrix, which correspond to the individual embeddings in this context.
To apply this concept to the convolution operation, the convolution weights, after unfolding, must form a rotation matrix. We define the dimensions of this matrix as $n=(channels_{in}\cdot kernel_{width}\cdot kernel_{height})$ and $n^{\prime}=channels_{out}$ respectively. However, the condition for orthogonal columns is not always satisfied as there will be instances where $n>n^{\prime}$. For these cases we use the norm-preserving transformation $\bm{z}=\bm{W^{T}}\bm{x}\cdot\frac{\lVert\bm{x}\rVert}{\lVert\bm{W^{T}}\bm{x}\rVert}$ where $\bm{W}\in\mathbb{R}^{(n^{\prime}\cdot n)}$. This formulation allows us to utilize existing efficient implementations of the convolution operation by directly parameterizing the kernel weights before passing them into the convolutional layer. Finally, we formalize the new Lorentz Convolution as:
$\bm{out}=\text{LorentzBoost(TanhScaling(RotationConvolution(}\bm{x}\textbf{)))}$)) (5)
where TanhRescaling is the operation described in Eq.3 and RotationConvolution is a normal convolution parameterized through the procedure in Algorithm 2, where Orthogonalize is a Cayley projection similar to (Helfrich et al., 2018) or the norm-preserving transformation above. We specifically use the Cayley projection because it always produces an orthonormal matrix with a positive determinant, ensuring that the rotated point remains on the upper sheet of the hyperboloid and avoids being mapped to the lower sheet.
Algorithm 2 Lorentz Convolution Parametarization
$\bm{W}\in\mathbb{R}^{C_{in},C_{out},K_{width},K_{length}}$
procedure Adaptweight(W)
     if $K_{width}\cdot K_{length}\cdot C_{in}>C_{out}$ then
         return $\bm{W}$
     else
         $\bm{W}=\bm{W}.reshape(K_{width}\cdot K_{length}\cdot C_{in},C_{out})$
         $\bm{\hat{W}}_{core}=$ Orthogonalize ($\bm{W}$)
         return $\bm{\hat{W}}.reshape(C_{in},C_{out},K_{width},K_{length})$
     end if
end procedure
$\bm{\hat{W}}_{core}$ can also be learned directly on the SPD manifold similar to Dai et al. (2021). This definition of the convolution operation allows to use the existing efficient implementations of 2D convolutions, saving both memory and runtime.
Lorentz-Core Bottleneck Block
To build on the concept of hybrid hyperbolic encoders introduced by (Bdeir et al., 2024), we developed the Lorentz Core Bottleneck blocks for hyperbolic ResNet-based models. These blocks are similar to standard Euclidean bottleneck blocks but replace the internal 3x3 convolutional layer with our efficient convolutional layer as illustrated in figure 2. This design allows us to incorporate hyperbolic structuring of the embeddings within each block while retaining the flexibility and computational efficiency of Euclidean models. We interpret this integration as a form of hyperbolic bias, enabling ResNets to leverage hyperbolic representations without requiring fully hyperbolic modeling.
Euclidean 1x1 ConvManifold Switcher + ScaleHyperbolic 3x3 ConvManifold Switcher + ScaleEuclidean 1x1 ConvBN + ReLUHyper BN + ReLU
Figure 2: Lorentz-Core Bottleneck Block
4 Experiments
4.1 Hierarchical Metric Learning Problem
In their paper Kim et al. (2023) take on the problem of hierarchical clustering using an unsupervised hyperbolic loss regularizer they name HIER. This method relies on the use of hierarchical proxies as learnable ancestors of the embedded data points in hyperbolic space. Given a triplet of points ${xi,xj,xk}$ where $x_{i}$ and $x_{j}$ are determined to be related by a reciprocal nearest neighbor measure, and $x_{k}$ is an unrelated point, the HIER loss regularizer is then calculated as
$\begin{split}\mathcal{L}_{\textrm{HIER}}(t)&=[D_{B}(x_{i},\rho_{ij})-D_{B}(x_{i},\rho_{ijk})+\delta]_{+}\\ &+[D_{B}(x_{j},\rho_{ij})-D_{B}(x_{j},\rho_{ijk})+\delta]_{+}\\ &+[D_{B}(x_{k},\rho_{ijk})-D_{B}(x_{k},\rho_{ij})+\delta]_{+},\end{split}$ (6)
where $D_{B}$ denotes the hyperbolic distance on the Poincaré ball, and $\rho_{ij}$ is the most likely least common ancestor of points $x_{i}$ and $x_{j}$. This encourages a smaller hyperbolic distance between $x_{i}$, $x_{j}$, and $\rho_{ij}$, and a larger distance with $\rho_{ijk}$. The opposite signal is then applied in the case of $x_{k}$, the irrelevant data point. Kim et al. (2023) show substantial performance uplifts for the HIER loss when applied to a variety of network architectures.
In the following experiment, we extend HIER to the Lorentz model (LHIER) and compare against the results provided by Kim et al. (2023).
Experimental Setup
We follow the experimental setup in Kim et al. (2023) and rely on four main datasets: CUB-200-2011 (CUB)(Welinder et al., 2010), Cars-196 (Cars)(Krause et al., 2013), Stanford Online Product (SOP)(Song et al., 2016), and In-shop Clothes Retrieval (InShop)(Liu et al., 2016). Performance is measured using Recall@k which is the fraction of queries with one or more relevant samples in their k-nearest neighbors. Additionally, all model backbones are pre-trained on Imagenet to ensure fair comparisons with the previous work.
Moving to Lorentz Space
To adapt the HIER model to the hyperboloid we first replace the Euclidean linear layer with a Lorentzian linear layer in the model neck and implement our max distance scaling operation after. We then set the hierarchical proxies as learnable hyperbolic parameters and optimize them directly on the manifold using our Lorentzian AdamW. Finally, we change the Poincaré distance to the Lorentz distance for the LHIER loss and set the hierarchical proxies to be scaled beforehand. We continue to use FP16 precision during curvature learning to evaluate the stability of the new optimization scheme. Our experiments also include both CNN-based and transformer models to assess whether differences in architecture affect the performance of the hyperbolic heads.
Results
As shown in table 1, our HIER+ model manages to improve the performance of mainly the CNN-based models with improvements ranging in recall@1 from 1-2.3%. In contrast, transformer models exhibit marginal gains in most cases and even show slight declines in other scenarios. This indicates that the proposed components do not better fully exploit the advantages of transformer architectures compared to other hyperbolic methods. The reasons may include architectural limitations or insufficient hyperparameter tuning. Notably, Kim et al. (2023) employed an extensive set of hyperparameters for model training, which we kept at default values to ensure fair comparisons and avoid costly hyperparameter searches.
Table 1: Performance of metric learning methods on the four datasets as provided by (Kim et al., 2023). All architecture backbones are pretrained and tested with the new LHIER loss. Superscripts denote their embedding dimensions and $\dagger$ indicates models using larger input images. As in (Kim et al., 2023), network architectures are abbreviated as, R–ResNet50 (He et al., 2016), B–Inception with BatchNorm (Ioffe & Szegedy, 2015), De–DeiT (Touvron et al., 2021), DN–DINO (Caron et al., 2021) and V–ViT (Dosovitskiy et al., 2021)
Methods Arch. CUB Cars SOP
R@1 R@2 R@4 R@1 R@2 R@4 R@1 R@10 R@100
Backbone architecture: CNN
NSoftmax (Zhai & Wu, 2018) R128 56.5 69.6 79.9 81.6 88.7 93.4 75.2 88.7 95.2
MIC (Roth et al., 2019) R128 66.1 76.8 85.6 82.6 89.1 93.2 77.2 89.4 94.6
XBM (Wang et al., 2020) R128 - - - - - - 80.6 91.6 96.2
XBM (Wang et al., 2020) B512 65.8 75.9 84.0 82.0 88.7 93.1 79.5 90.8 96.1
HTL (Ge et al., 2018) B512 57.1 68.8 78.7 81.4 88.0 92.7 74.8 88.3 94.8
MS (Wang et al., 2019) B512 65.7 77.0 86.3 84.1 90.4 94.0 78.2 90.5 96.0
SoftTriple (Qian et al., 2019) B512 65.4 76.4 84.5 84.5 90.7 94.5 78.6 86.6 91.8
PA (Kim et al., 2020) B512 68.4 79.2 86.8 86.1 91.7 95.0 79.1 90.8 96.2
NSoftmax (Zhai & Wu, 2018) R512 61.3 73.9 83.5 84.2 90.4 94.4 78.2 90.6 96.2
ProxyNCA++ (Teh et al., 2020) R512 69.0 79.8 87.3 86.5 92.5 95.7 80.7 92.0 96.7
Hyp (Ermolov et al., 2022) R512 65.5 76.2 84.9 81.9 88.8 93.1 79.9 91.5 96.5
HIER (Kim et al., 2023) R512 70.1 79.4 86.9 88.2 93.0 95.6 80.2 91.5 96.6
LHIER (ours) R512 72.4 81.5 88.4 89.1 93.5 96.1 81.3 92.1 96.8
Backbone architecture: ViT
[…]
Hyp (Ermolov et al., 2022) De128 74.7 84.5 90.1 82.1 89.1 93.4 83.0 93.4 97.5
HIER (Kim et al., 2023) De128 75.2 84.2 90.0 85.1 91.1 95.1 82.5 92.7 97.0
LHIER (ours) De128 75.5 84.7 90.6 85.4 91.7 95.6 82.7 93.5 97.4
Hyp (Ermolov et al., 2022) DN128 78.3 86.0 91.2 86.0 91.9 95.2 84.6 94.1 97.7
HIER (Kim et al., 2023) DN128 78.5 86.7 91.5 88.4 93.3 95.9 84.9 94.2 97.5
LHIER (ours) DN128 78.8 87.0 91.9 88.9 93.2 96.4 85.1 94.9 98.2
Hyp (Ermolov et al., 2022) V128 84.0 90.2 94.2 82.7 89.7 93.9 85.5 94.9 98.1
HIER  (Kim et al., 2023) V128 84.2 90.1 93.7 86.4 91.9 95.1 85.6 94.6 97.8
LHIER (ours) V128 84.6 90.2 93.9 86.7 92.0 95.3 85.9 95.0 98.0
[…]
DeiT-S (Touvron et al., 2021) De384 70.6 81.3 88.7 52.8 65.1 76.2 58.3 73.9 85.9
Hyp (Ermolov et al., 2022) De384 77.8 86.6 91.9 86.4 92.2 95.5 83.3 93.5 97.4
HIER (Kim et al., 2023) De384 78.7 86.8 92.0 88.9 93.9 96.6 83.0 93.1 97.2
LHIER (ours) De384 78.3 86.2 91.8 88.7 93.4 96.4 82.8 92.9 96.9
DINO (Caron et al., 2021) DN384 70.8 81.1 88.8 42.9 53.9 64.2 63.4 78.1 88.3
Hyp (Ermolov et al., 2022) DN384 80.9 87.6 92.4 89.2 94.1 96.7 85.1 94.4 97.8
HIER  (Kim et al., 2023) DN384 81.1 88.2 93.3 91.3 95.2 97.1 85.7 94.6 97.8
LHIER (ours) DN384 81.3 88.4 93.3 91.5 95.1 97.6 85.9 95.0 98.1
ViT-S (Dosovitskiy et al., 2021) V384 83.1 90.4 94.4 47.8 60.2 72.2 62.1 77.7 89.0
Hyp (Ermolov et al., 2022) V384 85.6 91.4 94.8 86.5 92.1 95.3 85.9 94.9 98.1
HIER  (Kim et al., 2023) V384 85.7 91.3 94.4 88.3 93.2 96.1 86.1 95.0 98.0
LHIER (ours) V384 86.2 92.1 95.2 88.6 93.6 96.3 86.7 95.3 98.3
4.2 Standard Classification Problem
Experimental Setup
We follow the experimental setup and hyperparameters of Bdeir et al. (2024) and rely on three main datasets: CIFAR10, CIFAR100, and Mini-Imagenet. We denote HCNN+ as the fully hyperbolic model from Bdeir et al. (2024), updated with efficient convolution implementations and our proposed scaling function. This experiment aims to boost model efficiency and address the performance inconsistencies observed in the original version.
For ResNet-18, we employ HECNN+, following the structure outlined in Bdeir et al. (2024) with alternating Euclidean and hyperbolic blocks, and integrate our efficient convolution operation. While we do not expect similar efficiency gains to those seen in fully hyperbolic models, our focus here is on enhancing scalability. In the case of ResNet-50, we adapt HECNN+ by replacing all blocks with the Lorentz-Core bottleneck block, where we anticipate the most significant efficiency improvements due to the larger model size. Our goal is to evaluate if the Lorentz-Core block can maintain strong performance under these conditions.
For both ResNet-18 and ResNet-50, we use Riemannian SGD with our improved learning scheme and curvature learning. We also decouple the encoder and decoder manifolds, allowing each to independently learn its own curvature for enhanced model flexibility.
Results
For the ResNet-18 experiments, Table 2 shows that the new architectures achieve comparable or better results across almost all cases. The smallest improvements are observed with the hybrid models, likely due to the reduced impact of the hyperbolic components compared to a fully hyperbolic model. However, we see a significant performance gap between the fully hyperbolic models, where our proposed architecture now matches the performance of the hybrid encoders. We hypothesize that the improved scaling helps mitigate the previous performance inconsistencies, with the whole model being two times more memory efficient.
Table 2: Classification accuracy (%) of ResNet-18 models. We estimate the mean and standard deviation from five runs. The best performance is highlighted in bold (higher is better).
CIFAR-10 CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.26)$ $(\delta_{rel}=0.23)$ For Cifar100 $(\delta_{rel}=0.20)$ Euclidean (He et al., 2016) ${95.14_{\pm 0.12}}$ $77.72_{\pm 0.15}$ 1.2GB - 12s $65.19_{\pm 0.12}$ Hybrid Poincaré (Guo et al., 2022) $95.04_{\pm 0.13}$ $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ Hybrid Lorentz (Bdeir et al., 2024) ${94.98_{\pm 0.12}}$ ${78.03_{\pm 0.21}}$ ${65.63_{\pm 0.10}}$ Poincaré ResNet (van Spengler et al., 2023) $94.51_{\pm 0.15}$ $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ HECNN Lorentz (Bdeir et al., 2024) ${95.16_{\pm 0.11}}$ ${78.76_{\pm 0.24}}$ 4.3GB - 100s ${65.96_{\pm 0.18}}$ HECNN+ (ours) ${95.15_{\pm 0.07}}$ ${78.80_{\pm 0.12}}$ 3GB - 80s ${65.98_{\pm 0.11}}$ HCNN Lorentz (Bdeir et al., 2024) ${95.15_{\pm 0.08}}$ ${78.07_{\pm 0.17}}$ 10GB - 175s ${65.71_{\pm 0.13}}$ HCNN+ (ours) ${95.17_{\pm 0.09}}$ $\bm{78.81_{\pm 0.19}}$ 5GB - 140s $\bm{66.12_{\pm 0.14}}$
In the ResNet-50 experiments in Table 3, we observe that HECNN+ significantly outperforms the Euclidean model across all datasets, including Tiny-ImageNet, where other models typically begin to show a drop in accuracy. This improvement is likely due to the tighter integration of hyperbolic components and the enhanced scaling, which helps manage the challenges of higher-dimensional embeddings.
We evaluate the impact of our efficient convolution and Lorentz-Core block in Table 3. We see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime. We attribute this improvement to the efficient closed-source convolution operations we can now leverage. Similar gains are also observed in generation tasks and classification with smaller models. However, there remains significant room for further optimization when compared to the Euclidean baseline. A key remaining performance bottleneck identified is the batch normalization step, which contributes approximately $60\%$ of the runtime and around $30\%$ of the memory usage. The next step would be to factorize the extensive parallel transports and tangent mappings involved in this operation, potentially alleviating the associated overhead.
Table 3: Classification accuracy (%) of ResNet-50 models. The best performance is highlighted in bold (higher is better).
CIFAR-10 CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.26)$ $(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ Euclidean (He et al., 2016) 95.1495.14{95.14}95.14 $78.52$ 4.5GB - 30s $66.23$ Hybrid Lorentz (Bdeir et al., 2024) 95.3895.38{95.38}95.38 ${79.35}$ 66.0166.01{66.01}66.01 HECNN (Bdeir et al., 2024) 95.4295.42{95.42}95.42 79.8379.83{79.83}79.83 15.6GB - 300s $66.30$ HCNN+ w Bottleneck Conv. (ours) $\bm{95.46}$ $\bm{80.86}$ 8.1GB - 100s $\bm{67.18}$
4.3 VAE Image Generation
Experimental Setup
Previous research has highlighted the ability of hyperbolic neural networks to perform better when using lower dimensional embeddings due to the more expressive space (Nagano et al., 2019; Mathieu et al., 2019; Ovinnikov, 2019; Hsu et al., 2020). A natural extension to this would be implementing the models for VAEs which rely on smaller latent embedding dimensions to encode the inputs. Bdeir et al. (2024) perform this study on image generation and reconstruction for their fully hyperbolic models. In the following section, we reproduce the experimental setup and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers. We also use curvature learning with our adjusted Riemannian SGD learning scheme.
Results
We see in table 4 that our implementation of the fully hyperbolic VAE achieves better performance on both datasets. It should also be noted that this is achieved with around 2.5x less memory and at 3x greater training speed. This helps demonstrate both the effectiveness of our proposed curvature learning process and our efficient model components.
Table 4: Reconstruction and generation FID of manifold VAEs across five runs (lower is better).
CIFAR-100 CelebA Rec. FID Gen. FID Rec. FID Gen. FID Euclidean $63.81_{\pm 0.47}$ $103.54_{\pm 0.84}$ $54.80_{\pm 0.29}$ $79.25_{\pm 0.89}$ Hybrid Poincaré (Mathieu et al., 2019) $62.64_{\pm 0.43}$ $98.19_{\pm 0.57}$ $54.62_{\pm 0.61}$ $81.30_{\pm 0.56}$ Hybrid Lorentz (Nagano et al., 2019) $62.14_{\pm 0.35}$ $98.34_{\pm 0.62}$ $54.64_{\pm 0.34}$ $82.78_{\pm 0.93}$ HCNN Lorentz (Bdeir et al., 2024) $61.44_{\pm 0.64}$ $100.27_{\pm 0.84}$ $54.17_{\pm 0.66}$ $78.11_{\pm 0.95}$ HCNN+ Lorentz (Ours) $57.69_{\pm 0.52}$ $98.14_{\pm 0.44}$ $52.73_{\pm 0.27}$ $77.98_{\pm 0.32}$
5 Conclusion
In our work, we present many new components and schemas for the use of hyperbolic deep learning in hyperbolic vision. We test these components in three vision tasks and prove the potential of these new components even in float16 conditions.
However, there is still significant room for improvement. Further optimizations to the batch normalization layers could enhance the efficiency of hyperbolic models. Additionally, a key challenge remains with the hyperbolic linear layers when reducing dimensionality. Currently, we match norms to simulate a rotation operation, but we encourage exploring alternative approaches that align more naturally with the mathematical properties of the manifold.
Appendix A Appendix
A.1 Operations in hyperbolic geometry
Parallel Transport
A parallel transport operation $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)$ describes the mapping of a vector on the manifold $\bm{v}$ from the tangent space of $\bm{x}\in\mathbb{L}$ to the tangent space of $\bm{y}\in\mathbb{L}$. This operation is given as $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)=\bm{v}+\frac{\langle\bm{y},\bm{v}\rangle_{\mathbb{L}}}{K-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}(\bm{x}+\bm{y}).$
Lorentzian Centroid (Law et al., 2019)
Also denoted as $\bm{\mu}_{\mathbb{L}}$, is the weighted centroid between points on the manifold based on the Lorentzian square distance. Given the weights $\bm{\nu}$, $\bm{\mu}=\frac{\sum_{i=1}^{m}\nu_{i}\bm{x}_{i}}{\sqrt{1/K}\left|||\sum_{i=1}^{m}\nu_{i}\bm{x}_{i}||_{\mathcal{L}}\right|}.$
Lorentz Transformations
In the Lorentz model, linear transformations preserving the structure of spacetime are termed Lorentz transformations. A matrix $\mathrm{\textbf{A}}^{(n+1)\times(n+1)}$ is defined as a Lorentz transformation if it provides a linear mapping from $\mathbb{R}^{n+1}$ to $\mathbb{R}^{n+1}$ that preserves the inner product, i.e., $\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{y}\rangle_{\mathbb{L}}=\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$ for all $\bm{x},\bm{y}\in\mathbb{R}^{n+1}$. The collection of these matrices forms an orthogonal group, denoted $\bm{O}(1,n)$, which is commonly referred to as the Lorentz group.
In this model, we restrict attention to transformations that preserve the positive time orientation, operating within the upper sheet of the two-sheeted hyperboloid. Accordingly, the transformations we consider lie within the positive Lorentz group, denoted $\bm{O}^{+}(1,n)={\mathrm{\textbf{A}}\in\bm{O}(1,n):a_{11}>0}$, ensuring preservation of the time component sign $x_{t}$ for any $\bm{x}\in\mathbb{L}^{n}_{K}$. Specifically, in this context, Lorentz transformations satisfy the relation
$\bm{O}^{+}(1,n)={\mathrm{\textbf{A}}\in\mathbb{R}^{(n+1)\times(n+1)}|\forall\bm{x}\in\mathbb{L}^{n}_{K}:\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{x}\rangle_{\mathbb{L}}=-\frac{1}{K},(\mathrm{\textbf{A}}\bm{x})_{0}>0)}.$ (7)
Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ Moretti (2002). The rotation matrix R is designed to rotate points around the time axis and is defined as
$\mathrm{\textbf{R}}=\left[\begin{array}[]{cc}1&\bm{0}^{T}\\ \bm{0}&\tilde{\mathrm{\textbf{R}}}\end{array}\right],$ (8)
where $\bm{0}$ represents a zero vector, $\tilde{\mathrm{\textbf{R}}}$ satisfies $\tilde{\mathrm{\textbf{R}}}^{T}\tilde{\mathrm{\textbf{R}}}=\mathrm{\textbf{I}}$, and $\det(\tilde{\mathrm{\textbf{R}}})=1$. This structure shows that Lorentz rotations on the upper hyperboloid sheet belong to a special orthogonal subgroup, $\bm{SO}^{+}(1,n)$, which preserves orientation, with $\tilde{\mathrm{\textbf{R}}}\in\bm{SO}(n)$.
In contrast, the Lorentz boost applies shifts along spatial axes given a velocity vector $\bm{v}\in\mathbb{R}^{n}$ with $||\bm{v}||<1$, without altering the time axis.
$\mathrm{\textbf{B}}=\left[\begin{array}[]{cc}\gamma&-\gamma\bm{v}^{T}\\ -\gamma\bm{v}&\mathrm{\textbf{I}}+\frac{\gamma^{2}}{1+\gamma}\bm{v}\bm{v}^{T}\end{array}\right],$ (9)
with $\gamma=\frac{1}{\sqrt{1-||\bm{v}||^{2}}}$. However, this can also be any operation that scales the norms of the space values without changing the vector orientation.
A.2 Scaling Lorentzian Vectors
Tanh Scaling
We show the output of the tanh scaling function in Figure 3. By changing the transformation parameters we are able to fine-tune the maximum output and the slope to match our desired function response.
Figure 3: The output of the proposed flexible tanh function. Here the maximum value m is set to 9.1 in the vanilla version with an alternate value of m=18 and the slope s is set to 2.6 with an alternate value of 3.5
Hyperbolic Scaling
We isolate the transformation of the $\exp^{K}_{\overline{\bm{0}}}\left(y\right)$ operation on the space values of $y$ as:
$\bm{x}_{s}=\sqrt{K}\times sinh(\frac{\lVert\bm{y}\rVert_{\mathbb{L}}}{\sqrt{K}})\frac{\bm{y}}{\lVert\bm{y}\rVert_{\mathbb{L}}}$ (10)
where $\bm{y}\in\mathbb{R}^{d}=\log^{K}_{\overline{\bm{0}}}(\bm{x})$. However, at the tangent plane of the origin, the first element $\bm{y}_{0}$ becomes 0. As such $\lVert\bm{y}\rVert_{\mathbb{L}}=\lVert\bm{y}\rVert_{\mathbb{E}}=\sum_{i=2}^{d}\bm{y_{i}^{2}}$. This gives us:
$\bm{x}_{s}=\sqrt{K}\times sinh(\frac{\lVert\bm{y}\rVert_{\mathbb{E}}}{\sqrt{K}})\frac{\bm{y}}{\lVert\bm{y}\rVert_{\mathbb{E}}}$ (11)
We can now scale the norm of the Euclidean vector $\bm{y}$ bay a value $a$ and find the equivalent value for the hyperbolic space elements:
$a_{\mathbb{L}}=\frac{\bm{x}_{s_{rescaled}}}{\bm{x}_{s}}=\frac{sinh(\frac{a\times\lVert\bm{y}\rVert_{\mathbb{E}}}{\sqrt{K}})}{sinh(\frac{\lVert\bm{y}\rVert_{\mathbb{E}}}{\sqrt{K}})}=\frac{e^{\frac{a\times\lVert\bm{y}\rVert_{\mathbb{E}}}{\sqrt{K}}}-e^{\frac{-a\times\lVert\bm{y}\rVert_{\mathbb{E}}}{\sqrt{K}}}}{e^{\frac{\lVert\bm{y}\rVert_{\mathbb{E}}}{\sqrt{K}}}-e^{\frac{-\lVert\bm{y}\rVert_{\mathbb{E}}}{\sqrt{K}}}}$ (12)
Additionally, we know that the hyperbolic distance from the origin of the manifold to any point is equal to the norm of the projected vector onto the tangent plane. Supposing that we want $a\times D(\bm{x},\overline{\bm{0}})^{K}=D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}$, we get the final equation:
$\bm{x}_{s_{rescaled}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}$ (13)
A.3 Ablations
We test the effect of individual model components in table 5. Each subsequent model involves the default architecture presented in the experimental setup minus the mentioned component. As we can see, the best results are achieved when all the architectural components are included. In the case of attempting to learn the curvature without the proposed optimizer schema, the model breaks completely down due to excessive numerical inaccuracies. One other benefit that we find from learning the curvature is quicker convergence. The model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.
Table 5: Ablation experiments for Resnet-50 models on Cifar-100.
CIFAR-100 HCNN+ - Default $\bm{80.86}$ HCNN+ - fixed curve $79.6$ HCNN+ - no scaling $80.13$ HCNN+ - no optim scheme $NaaN$

### v3

@@
-Optimizing Learning for Robust Hyperbolic Deep Learning in Computer Vision
+Robust Hyperbolic Learning with Curvature-Aware Optimization
 Abstract
-Hyperbolic deep learning has become a growing research direction in computer vision for the unique properties afforded by the alternate embedding space.
+Hyperbolic deep learning has become a growing research direction in computer vision due to the unique properties afforded by the alternate embedding space.
 The negative curvature and exponentially growing distance metric provide a natural framework for capturing hierarchical relationships between datapoints and allowing for finer separability between their embeddings.
-However, these methods are still computationally expensive and prone to instability, especially when attempting to learn the negative curvature that best suits the task and the data.
-Current Riemannian optimizers do not account for changes in the manifold which greatly harms performance and forces lower learning rates to minimize projection errors.
-Our paper focuses on curvature learning by introducing an improved schema for popular learning algorithms and providing a novel normalization approach to constrain embeddings within the variable representative radius of the manifold.
-Additionally, we introduce a novel formulation for Riemannian AdamW, and alternative hybrid encoder techniques and foundational formulations for current convolutional hyperbolic operations, greatly reducing the computational penalty of the hyperbolic embedding space.
-Our approach demonstrates consistent performance improvements across both direct classification and hierarchical metric learning tasks while allowing for larger hyperbolic models.
+However, current hyperbolic learning approaches are still prone to overfitting, computationally expensive, and prone to instability, especially when attempting to learn the manifold curvature to adapt to tasks and different datasets.
+To address these issues, our paper presents a derivation for Riemannian AdamW that helps increase hyperbolic generalization ability.
+For improved stability, we introduce a novel fine-tunable hyperbolic scaling approach to constrain hyperbolic embeddings and reduce approximation errors.
+Using this along with our curvature-aware learning schema for Lorentzian Optimizers enables the combination of curvature and non-trivialized hyperbolic parameter learning.
+Our approach demonstrates consistent performance improvements across Computer Vision, EEG classification, and hierarchical metric learning tasks achieving state-of-the-art results in two domains and drastically reducing runtime.
+Machine Learning, ICML
 1 Introduction
-With the recent rise in the use of hyperbolic manifolds for deep representation learning, there is a growing need for efficient, flexible components that can fully exploit these spaces without sacrificing stability.
-This has led researchers to focus on two main derivations of hyperbolic space: the Poincaré manifold and the hyperboloid.
-The Poincaré ball, equipped with a gyrovector space, supports various well-defined operations, including generalized vector addition and multiplication, but it suffers from significant stability issues.
-On the other hand, the hyperboloid, or Lorentz space, lacks these operations but offers much better operation stability, as demonstrated in the study by Mishne et al.
+Recently, hyperbolic manifolds have gained attention in deep learning for their ability to model hierarchical and tree-like data structures efficiently.
+Unlike Euclidean space, hyperbolic space has negative curvature, allowing it to represent exponentially growing distances.
+This makes these manifolds ideal for tasks like natural language processing, graph representation, and recommendation systems (Peng et al., 2022).
+By embedding data into hyperbolic space, models can capture complex relationships more effectively, often with fewer parameters (Ganea et al., 2018), leading to improved performance and interpretability.
+Hyperbolic geometry has also found applications in computer vision, where its ability to better separate embeddings in high-dimensional spaces has shown promise in improving tasks like image classification and segmentation(Bdeir et al., 2024; van Spengler et al., 2023; Khrulkov et al., 2020; Atigh et al., 2022), few-shot learning(Liu et al., 2020), and feature representation (Yang et al., 2023).
+Two main derivations of hyperbolic space have emerged as learning frameworks: the hyperboloid or the Lorentz space ($\mathbb{L}$) and the Poincaré manifold ($\mathbb{P}$).
+Typically, the hyperboloid, or Lorentz space, offers better operational stability, as demonstrated by Mishne et al.
+(2022) but lacks clear definitions for basic vector operations such as addition and subtraction.
+To bridge this gap, recent research has focused on defining Lorentzian versions of common deep learning operations, such as the feed-forward layer (Chen et al., 2022; Dai et al., 2021; Ganea et al., 2018), convolutional layer (Chen et al., 2022; Qu & Zou, 2023; Dai et al., 2021), and MLRs (Bdeir et al., 2024).
+However, the use of these hyperbolic operations comes with challenges.
+Computations in hyperbolic space are more complex and expensive due to the increased mathematical complexity and the lack of optimized CUDA implementations, which can slow down training and drastically increase memory requirements.
+Additionally, optimization on hyperbolic manifolds can be unstable, especially in low floating-point precision environments.
+This has required research to rely on parameter clamping techniques, which can cause non-smooth gradient updates, to remain within the accurate representation radius for manifolds Mishne et al.
+(2022); Guo et al.
 (2022).
-To address this gap, previous works have sought to provide Lorentzian definitions for common deep learning operations such as the feed-forward layer (Chen et al., 2022; Dai et al., 2021; Ganea et al., 2018), convolutional layer (Chen et al., 2022; Qu & Zou, 2023; Dai et al., 2021), and MLRs (Bdeir et al., 2024).
-This increased focus on hyperbolic modeling has led to its gradual integration into computer vision architectures, as detailed in the survey by Mettes et al.
-(2023).
-Specifically, the hyperboloid model has been employed as a sampling space for VAEs (Nagano et al., 2019), a decoder space for vision tasks in hybrid settings (Guo et al., 2022; Liu et al., 2020; Khrulkov et al., 2020; Qu & Zou, 2022), and ultimately for fully hyperbolic Lorentzian vision encoders (Bdeir et al., 2024) simultaneously with its Poincaré counterpart (van Spengler et al., 2023).
-This paper furthers the development of hyperbolic learning for vision tasks, specifically for the Lorentz manifold.
-Our primary focus is on the challenge of learning the manifold’s negative curvature.
-The driving principle behind this, is that the model embeddings may exhibit varying degrees of hyperbolicity depending on the innate hierarchies in the datapoints themselves, the problem task that is being considered, and the specific locations of hyperbolic operation integrations.
-To accomodate for this, we can adjust the embedding space’s hyperbolic metric to be less or more Euclidean which accounts for the modeling requirements.
-We also build on the idea of separate manifolds for separate main blocks in the architecture further increasing representative flexibility.
-We also recognize that despite recent advances, Lorentz models continue to struggle with issues of high computational costs.
-We attempt to isolate and alleviate the main factors leading to numerical inaccuracies and computational overhead overall, and more particularly when modeling data in higher-dimensional embedding spaces and when learning the curvatures.
-Our contributions can then be summed up as:
+Otherwise, performance either suffers or training becomes impossible due to the precision errors causing mathematically undefined operations.
+The instability is only exacerbated when we incorporate the negative curvature as a learned parameter since it directly affects the representational radius and requires the projection of vectors between these different curvatured manifolds.
+Finally, in data-scarce scenarios, the higher representational capacity makes the hyperbolic spaces more prone to overfitting which heavily impacts their generalization ability.
+This work addresses key challenges in the Lorentz model.
+To combat overfitting, we introduce an AdamW-based optimizer that enhances regularization and generalization in hyperbolic learning.
+For learning instability, we propose an optimization framework that stabilizes curvature parameter learning.
+We also present a smoother scaling function to replace weight clipping, ensuring hyperbolic vectors remain within the representational radius.
+To tackle computational complexity, we introduce an implementation trick that leverages efficient CUDA-based convolutions for hyperbolic learning.
+Improved stability not only boosts model performance but also enables lower-precision training, enhancing efficiency.
+Our contributions are four-fold:
   1. 1.
-    We propose a formulation for Riemannian AdamW and an alternative schema for Riemannian optimizers that accounts for manifold curvature learning.
+    We propose a formulation for Riemannian AdamW and an alternative schema for Riemannian optimizers that stabilize curvature learning for hyperbolic parameters.
   2. 2.
@@
   3. 3.
-    We provide a more efficient convolutional layer approach that is able to leverage the highly optimized existing implementations.
+    We present HCNN+, HyperMAtt, and LHEIR, an application of our proposed optimization scheme for the domains of classical image classification, EEG classification, and hierarchical metric learning respectively.
   4. 4.
-    We empirically show the effectiveness of combining these approaches using classical image classification tasks and hierarchical metric learning problems.
+    We empirically show the effectiveness of our proposed methods in three domains, hierarchical metric learning, Image, and EEG classification, where we achieve state-of-the-art performance in two domains with a significant computational speed-up.
+Our code is publicly available at [redacted for double-blind submission].
 2 Related Work
-Hyperbolic Embeddings in Computer Vision
+Hyperbolic Embeddings in Deep Learning
 With the success of employing hyperbolic manifolds in NLP models (Zhu et al., 2021; Dhingra et al., 2018; Tifrea et al., 2018) hyperbolic embeddings have extended to the computer vision domain.
@@
 In comparison, our proposed optimization schema maintains the parameters on the manifold and optimizes them directly by performing the necessary operations to transition between the variable curvature spaces.
-Metric Learning
-Metric learning relies on the concept of structuring the distribution in the embedding space so that related data points are positioned closer together, while less related points are placed further apart.
-To facilitate this process, numerous studies have introduced additional loss functions that explicitly encourage this behavior.
-Contrastive losses, for instance, operate on pairs of data points and propose a penalty that is proportional to the distances between negative pairs and inversely proportional to the distance between positive pairs (Chopra et al., 2005).
-Triplet loss extends this idea by considering sets of three points: an anchor, a positive sample, and a negative sample (Wang et al., 2014).
-Instead of changing the distances between points absolutely, it ensures that the distance between the anchor and the positive sample is less than the distance between the anchor and the negative sample, plus a margin, thus enforcing a relational criterion.
-These approaches have also been adapted to hierarchical problem settings under hyperbolic manifolds (Yang et al., 2022; Kim et al., 2023).
-Notably, Kim et al.
-(2023) developed a method for learning continuous hierarchical representations using a deep learning, data-mining-like approach that relies on the innate relationships of the embeddings rather than their labels.
-They employ a proxy-based method that models the data on the Poincaré ball, facilitating a more natural extension to hierarchical tasks.
-Building on this, we extend the approach by modeling the loss function in the Lorentz manifold and incorporating a learnable curvature to better handle data with varying levels of hierarchy.
 3 Methodology
-Hyperbolic embeddings offer significant advantages for computer vision tasks that benefit from hierarchical structure modeling, but their effectiveness is hindered by practical challenges, including mathematical inconsistencies in curvature learning, the absence of a Riemannian AdamW optimizer, stability issues with float32 precision, and high computational costs from inefficient convolutions.
-To address these issues, the following sections introduce four key advancements: a more stable curvature learning method, a Riemannian AdamW optimizer, a normalization scheme for improved float32 precision stability, and efficient hyperbolic convolutions to reduce memory and computational costs.
-3.1 Background
+3.1 The Lorentz Space
 The hyperbolic space is a Riemannian manifold with a constant negative sectional curvature $c<0$.
 There are many conformal models of hyperbolic space but we focus our work on the hyperboloid, or Lorentz manifold.
-The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}\in\mathbb{R}^{n+1}~{}|~{}\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=\frac{-1}{K},~{}x_{t}>0\}$ where $\frac{-1}{K}=c$, and with the Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\mathrm{diag}(-1,1,\dots,1)$.
-This models the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}=[\sqrt{K},0,\cdots,0]^{T}$.
-We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector, the space dimension $\bm{x_{s}}$.
-The Lorentzian inner product then becomes $\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}=\bm{x}^{T}\mathrm{diag}(-1,1,\cdots,1)\bm{y}.$ We now define the common hyperbolic operations in the Lorentz space.
+The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}=[x_{t},\bm{x_{s}}]\in\mathbb{R}^{n+1}~{}|~{}\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=\frac{-1}{K},~{}x_{t}>0\}$ where $\frac{-1}{K}=c$ and the Lorentzian inner product as Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}.$ It then follows that ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ which is an important characteristic to note for the stability issues presented later on.
+The formulation $\mathcal{L}^{n}$ then presents the Lorentz manifold as the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}=[\sqrt{K},0,\cdots,0]^{T}$.
+We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector, as the space dimension $\bm{x_{s}}$.
+Below are the basic operations presented by the manifold with more complex operations provided in Appendix A.
 Distance
-Distance in hyperbolic space is the magnitude of the geodesic forming the shortest path between two points.
-Let $\bm{x},\bm{y}\in\mathbb{L}^{n}_{K}$, the distance between them is given by $d_{\mathbb{L}}(\bm{x},\bm{y})=\sqrt{K}\cosh^{-1}(\frac{-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}{K})$.
-We also define the square distance by Law et al.
-(2019) as $d^{2}_{\mathbb{L}}(\bm{x},\bm{y})=||\bm{x}-\bm{y}||_{\mathbb{L}}^{2}=-2K-2\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$.
+Distance in hyperbolic space is the magnitude of the geodesic forming the shortest path between two points. Let $\bm{x},\bm{y}\in\mathbb{L}^{n}_{K}$, the distance between them is given by
+$d_{\mathbb{L}}(\bm{x},\bm{y})=\sqrt{K}\cosh^{-1}(\frac{-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}{K})$ (1)
+We also define the square distance by Law et al. (2019) as $d^{2}_{\mathbb{L}}(\bm{x},\bm{y})=||\bm{x}-\bm{y}||_{\mathbb{L}}^{2}=-2K-2\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$.
 Exponential and Logarithmic Maps
-Seeing as the Lorentz space is a Riemannian manifold, it is locally Euclidean.
+Since the Lorentz space is a Riemannian manifold, it is locally Euclidean.
 This can best be described through the tangent space $\mathbb{T}_{\bm{x}}\mathcal{M}$, a first-order approximation of the manifold at a given point $\bm{x}$.
 The exponential map, $\exp^{K}_{\bm{x}}\left(\bm{z}\right):\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}\rightarrow\mathbb{L}^{n}_{K}$ is then the operation that maps a tangent vector $\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}$ onto the manifold through $\exp_{\bm{x}}^{K}(\bm{z})=\cosh(\alpha)\bm{x}+\sinh(\alpha)\frac{\bm{z}}{\alpha},\text{ with }\alpha=\sqrt{1/K}||\bm{z}||_{\mathbb{L}},~{}~{}||\bm{z}||_{\mathbb{L}}=\sqrt{{\langle\bm{z},\bm{z}\rangle}_{\mathbb{L}}}.$ The logarithmic map is the inverse of this mapping and can be described as $\log_{\bm{x}}^{K}(\bm{y})=\frac{\cosh^{-1}(\beta)}{\sqrt{\beta^{2}-1}}\cdot(\bm{y}-\beta\bm{x}),\text{ with }\beta=-\frac{1}{K}\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}.$
-3.2 Riemannian Optimization
-On the Stability of Curvature Learning
-In the hyperbolic learning library GeoOpt, Kochurov et al.
-(2020) introduce the curvature of hyperbolic space as a learnable parameter.
-However, no subsequent work has fully leveraged this feature, and our empirical testing reveals that this approach often leads to instability and performance decline.
-We attribute these issues to the simplistic implementation of the curvature updates, which neglects to update the hyperbolic operations after updating the curvature.
-In the current Riemannian optimization approach, Euclidean gradients and their momentums are projected onto the Riemannian manifold and then used to update the hyperbolic parameters.
-These operations inherently depend on the curvature of the space.
-However, when the curvature is adjusted during learning, the projections, gradients, and hyperbolic parameters can become inconsistent, as they mix the projection and update operations from the new curvature with vectors defined under the previous curvature.
-To illustrate this, Figure 1 shows two manifolds with different curvatures.
+3.2 Riemannian Optimization Schema
+Background
+In the GeoOpt hyperbolic learning library, Kochurov et al.
+(2020) introduces the curvature $c$ of hyperbolic space as a learnable parameter.
+However, no subsequent work has fully taken advantage of this feature despite the authors presenting ready implementations of hyperbolic operations that account for this curvature.
+Our empirical testing also reveals that naively learning this parameter often leads to instability and performance decline.
+We attribute these issues to the simplistic curvature updates in Riemannian optimizers, which neglect to adjust hyperbolic operations, weights, and gradients after updating the curvature.
+In the current Riemannian optimization approach, Euclidean gradients and momentums are projected onto the manifold to update hyperbolic parameters.
+However, since optimization depends on curvature, adjusting it during learning makes previous projections, gradients, and parameters inconsistent with the new geometry.
+Essentially, the current schema applies updates using the new curvature on vectors defined by the old curvature, leading to mismatches.
+To illustrate this, LABEL:fig:curve_inconsistencies shows two manifolds with different curvatures.
 It is evident that points defined on the first manifold do not lie on the second, and the tangent planes of the first manifold are not aligned with those of the second.
 Thus, using the new curvature operations on points from the old model is not mathematically sound practice.
-This inconsistency introduces instability, undermining the reliability of curvature learning and often leading to invalid values during the parameter updates.
-To address this, we propose a method that tracks both the old and new curvature values throughout updates.
-Specifically, all updates are applied under the old curvature, after which the hyperbolic vectors are mapped from the old manifold to the new one.
-While GeoOpt does not address the optimization challenges involved in learning manifold curvature, it does implement an ”N-stabilize step”.
+This inconsistency introduces instability, undermining the reliability of curvature learning and often leading to undefined values during the parameter updates.
+One example of this is the distance function Equation 1 which requires the bound ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ to hold.
+If the new curvature parametrization at time t $K_{t}>K_{t-1}$ then the input into the inverse hyperbolic cosine has a lower bound less than 1, leaving room for the operation to be mathematically undefined.
+While GeoOpt does not address this particular optimization challenge, it does implement an ”N-stabilize step”.
 This step aims to reduce instability by recalculating the time component of the hyperbolic parameters every N-steps to ensure they remain on the manifold.
-However, this approach fails to prevent invalid parameter values before the update steps and suffers from high inaccuracies due to the mathematically flawed parameter optimization.
-Moreover, the time re-calculation projection used in the stabilize step alters the relative magnitudes of the hyperbolic vectors and gradients, and even the directions of the momentums, which can degrade performance and lead to training instability.
-Instead, we recommend projecting points onto the tangent space at the old origin using the logarithmic map and then projecting back after the curvature update.
-Gradients and their momentum can be similarly projected or parallel transported to the tangent space.
-Parallel transport is chosen here because it preserves the vectors’ directions which is important for the optimization procedure.
-Additionally, when the curvature changes, the tangent space at the origin remains relatively stable, as shown in Figure 1.
-Moving onto the new manifold is then a simple exponential map and parallel transport back operation.
-This preserves the vector norms and their hyperbolic distances to the origin since $D(\exp^{K_{1}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{1}}=||y||=D(\exp^{K_{2}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{2}}$ where $\bm{y}\in\mathcal{T}_{\overline{\bm{0}}}\mathcal{M}$.
-No changes are needed for the tangent vector during the move between tangent spaces of different curvatures since these spaces are parallel, making the transition a simple translation along the time axis.
-(a) At a random point
-(b) At the origin
-Figure 1: Tangent planes of a hyperboloid with curvature -1 relative to another hyperboloid with curvature -0.7.
-Tangential properties between manifolds are better respected at the origin where tangents remain parallel.
+However, this approach is done after the update step and does not prevent invalid parameter values during optimization.
+Curvature Aware Optimization
+To address this, we propose a method that tracks both the old and new curvature parameter values $\bm{\theta}_{K}$ throughout updates.
+Specifically, we propose a simple approach where all updates are first applied under the old curvature, then the new curvature is derived and finally, the hyperbolic vectors are mapped from the old manifold to the new one.
+The previously proposed mapping technique in the N-stabilize step alters the relative magnitudes of the hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ and gradients $\mathcal{G}$, as well as the directions of the momentums, which can degrade performance and lead to training instability.
+Instead, we recommend projecting points onto the tangent space at the origin of the old curvature $\overline{\bm{0}}_{t-1}$ using the logarithmic map and then projecting back after the curvature update.
+Gradients and their momentum can be similarly projected or parallel transported Appendix A) to the tangent space.
+The entire optimization schema can then be seen in Algorithm 1.
+Here, we choose to apply parallel transport, or any similar approximation of it, because it preserves the gradient vectors’ directions which is important for the optimization procedure.
+We also choose to use the tangent space at the origin as the mapping intermediate since it remains relatively stable when the curvature is updated, as shown in LABEL:fig:curve_inconsistencies.
+Finally, we prove that the vector norms and their hyperbolic distances to the origin are preserved since $D(\exp^{K_{1}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{1}}=||y||=D(\exp^{K_{2}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{2}}$ where $\bm{y}\in\mathcal{T}_{\overline{\bm{0}}}\mathcal{M}$.
 It is important to emphasize that our proposed optimization scheme is compatible with existing curvature learning and meta-learning methods.
 Rather than being an alternative, it serves as a guideline for updating manifold parameters during curvature changes, aimed at maintaining learning stability throughout the process.
-Riemannian AdamW Optimizer
-Recent works, particularly with transformers, commonly use the AdamW optimizer introduced by Loshchilov & Hutter (2019).
-As of current, there is no established Riemannian variant of this optimizer.
-We attempt to derive AdamW for the Lorentz manifold and argue a similar approach could be extended to the Poincaré ball.
-The main difference between AdamW and Adam is the direct weight regularization.
-This is more difficult to perform in the Lorentz space given the lack of an intuitive subtraction operation on the manifold.
-To resolve this, we reframe the parameter regularization in the original AdamW as a weighted centroid with the origin $O$
-$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda O$
+3.3 Riemannian AdamW Optimizer
+Background
+The AdamW optimizer was first introduced by Loshchilov & Hutter (2019) and relies on an improved application of the L2-regularization factor in the base Adam optimizer.
+L2-regularization works on the principle that networks with smaller weights tend to have better generalization performance than equal networks with higher weight values.
+Adam applied the L2-regularization during the gradient update step by incorporating it in the loss.
+However, Loshchilov & Hutter (2019) argue that this is inconsistent since the regularization effect is reduced by the magnitude of the gradient norms.
+Instead, they apply the weight decay directly during the parameter update step.
+AdamW is then shown to generalize better and lead to better convergence and has become a popular choice for many vision tasks.
+Given the above, we believe AdamW is significant for hyperbolic learning.
+Hyperbolic spaces, with their higher representational capacity, are prone to overfitting, especially under data scarcity during training (Gao et al., 2022).
+Thus, AdamW’s enhanced L2-regularization could improve hyperbolic models.
+Riemannian AdamW
+In the following, we derive AdamW for the Lorentz manifold and suggest its extension to the Poincaré ball.
+The key difference between AdamW and Riemannian Adam lies in direct weight regularization, which is challenging in Lorentz space due to the lack of an intuitive subtraction operation.
+To address this, we re-frame parameter regularization in AdamW as a weighted centroid with the origin $\overline{\bm{0}}$
+$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda\overline{\bm{0}}$
 where $\gamma$ is the learning rate and $\lambda$ is the weight decay value.
@@
 where $\mu_{\mathbb{L}}$ is the Lorentz centroid, and $\bm{\nu}$ are the weights. By removing the later gradient decay and introducing this operation, we adapt AdamW for use in the Lorentz space.
-Algorithm 1 Curvature Learning Aware Optimization Given parameters $\bm{\theta}=[\bm{\theta}_{euclid},\bm{\theta}_{\mathbb{L}},\bm{\theta}_{K}]$ and gradients denoted $\mathcal{G}$ procedure Optimizer Step($\bm{\theta}$)      for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do          Vanilla Riemannian Optimizer Step      end for      for $\bm{p}$ in $\bm{\theta}_{euclid}\cup\bm{\theta}_{K}$ do          Vanilla Eucldiean Optimizer Step      end for      MoveParameters() end procedure procedure Move Parameters      for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do          $\bm{\mathcal{G}}_{temp}=\mathrm{PT}^{K_{t-1}}_{{\bm{p}}\rightarrow{\bm{\overline{\bm{0}}}_{t-1}}}\left({\bm{\mathcal{G}}}\right)$          $\bm{z}=\log_{\bm{\overline{\bm{0}}}_{t-1}}^{K_{t-1}}(\bm{p})$          $\bm{p}=\exp_{\overline{\bm{0}}_{t}}^{K_{t}}(\bm{z})$          $\bm{\mathcal{G}}=\mathrm{PT}^{K_{t}}_{{\bm{\overline{\bm{0}}}_{t}}\rightarrow{\bm{p}}}\left({\bm{\mathcal{G}}_{temp}}\right)$      end for end procedure
-Maximum Distance Rescaling
+Algorithm 1 Curvature Learning Aware Optimization
+  Given parameters $\bm{\theta}=[\bm{\theta}_{euclid},\bm{\theta}_{\mathbb{L}},\bm{\theta}_{K}]$ and gradients denoted $\mathcal{G}$
+  function Optimizer Step
+     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
+        Vanilla Riemannian Optimizer Step
+     end for
+     for $\bm{p}$ in $\bm{\theta}_{euclid}\cup\bm{\theta}_{K}$ do
+        Vanilla Euclidean Optimizer Step
+     end for
+     MoveParameters()
+  end function
+  function Move Parameters
+     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
+        $\bm{\mathcal{G}}_{temp}=\mathrm{PT}^{K_{t-1}}_{{\bm{p}}\rightarrow{\bm{\overline{\bm{0}}}_{t-1}}}\left({\bm{\mathcal{G}}}\right)$
+        $\bm{z}=\log_{\bm{\overline{\bm{0}}}_{t-1}}^{K_{t-1}}(\bm{p})$
+        $\bm{p}=\exp_{\overline{\bm{0}}_{t}}^{K_{t}}(\bm{z})$
+        $\bm{\mathcal{G}}=\mathrm{PT}^{K_{t}}_{{\bm{\overline{\bm{0}}}_{t}}\rightarrow{\bm{p}}}\left({\bm{\mathcal{G}}_{temp}}\right)$
+     end for
+  end function
+3.4 Maximum Distance Rescaling
+Background
 […]
@@
 (2022).
+Flexible Scaling Function
 To address these issues, we introduce a modified scaling function, designed to provide finer control over both the maximum values and the slope of the curve.
@@
 To apply this to the hyperbolic vectors, we suggest performing the scaling on the tangent plane of the origin.
-However, this is an expensive operation to perform often, as such we derive in appendix A.2 the equivalent factorized form for the scaling of the space values:
+However, this is an expensive operation to perform repeatedly, as such we derive in Appendix C the equivalent factorized form for the scaling of the space values:
 ${\bm{x}_{s_{rescaled}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}}$ (4)
-where $D(\bm{x},\overline{\bm{0}})_{rescaled}^{k}$ is the scaled distances by plugging in $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Eq.2.
-The time component of the vector is then recalculated based on the norm of its spatial components.
-This gives us a complete scaling operation for the Lorentz space.
-We apply this tanh scaling when moving parameters across different manifolds.
-This includes transitions from the Euclidean space to the Lorentz space, as well as between Lorentz spaces of different curvatures.
-Additionally, we use this scaling after Lorentz Boosts and direct Lorentz concatenations (Qu & Zou, 2022).
-We also incorporate it following the variance-based rescaling in the batch normalization layer, since variance adjustments during this operation can push points outside the radius during the operation and often lead to invalid values.
-3.3 Towards Efficient Architectural Components
-Lorentz Convolutional Layer
-In their work, Bdeir et al.
-(2024) proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al.
-(2022).
-This approach ensured that the convolution outputs remained on the hyperboloid.
-However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions.
-To address this issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al.
-(2021), which decomposes the transformation into a Lorentz boost and a Lorentz rotation.
-Using this definition, we replace the matrix multiplication employed by Bdeir et al.
-(2024) for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost.
-Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
-A rotation is any operation that modifies the directions of transformed vectors while preserving their magnitudes. This can be achieved using a matrix $\bm{\hat{W}}\in\mathbb{S}$. Where
-$\mathbb{S}(n^{\prime},n)={\bm{M}\in\mathbb{R}^{(n^{\prime}\cdot n)}:\bm{M}^{T}\bm{M}=\bm{I}}$ (4)
-represents the Stiefel manifold.
-The condition $n\leq n^{\prime}$ ensures that the columns of the weight matrix are orthogonal, thereby preserving the norms along the rows of the input matrix, which correspond to the individual embeddings in this context.
-To apply this concept to the convolution operation, the convolution weights, after unfolding, must form a rotation matrix.
-We define the dimensions of this matrix as $n=(channels_{in}\cdot kernel_{width}\cdot kernel_{height})$ and $n^{\prime}=channels_{out}$ respectively.
-However, the condition for orthogonal columns is not always satisfied as there will be instances where $n>n^{\prime}$.
-For these cases we use the norm-preserving transformation $\bm{z}=\bm{W^{T}}\bm{x}\cdot\frac{\lVert\bm{x}\rVert}{\lVert\bm{W^{T}}\bm{x}\rVert}$ where $\bm{W}\in\mathbb{R}^{(n^{\prime}\cdot n)}$.
-This formulation allows us to utilize existing efficient implementations of the convolution operation by directly parameterizing the kernel weights before passing them into the convolutional layer.
-Finally, we formalize the new Lorentz Convolution as:
-$\bm{out}=\text{LorentzBoost(TanhScaling(RotationConvolution(}\bm{x}\textbf{)))}$)) (5)
-where TanhRescaling is the operation described in Eq.3 and RotationConvolution is a normal convolution parameterized through the procedure in Algorithm 2, where Orthogonalize is a Cayley projection similar to (Helfrich et al., 2018) or the norm-preserving transformation above.
-We specifically use the Cayley projection because it always produces an orthonormal matrix with a positive determinant, ensuring that the rotated point remains on the upper sheet of the hyperboloid and avoids being mapped to the lower sheet.
-Algorithm 2 Lorentz Convolution Parametarization
-$\bm{W}\in\mathbb{R}^{C_{in},C_{out},K_{width},K_{length}}$
-procedure Adaptweight(W)
-     if $K_{width}\cdot K_{length}\cdot C_{in}>C_{out}$ then
-         return $\bm{W}$
-     else
-         $\bm{W}=\bm{W}.reshape(K_{width}\cdot K_{length}\cdot C_{in},C_{out})$
-         $\bm{\hat{W}}_{core}=$ Orthogonalize ($\bm{W}$)
-         return $\bm{\hat{W}}.reshape(C_{in},C_{out},K_{width},K_{length})$
-     end if
-end procedure
-$\bm{\hat{W}}_{core}$ can also be learned directly on the SPD manifold similar to Dai et al.
-(2021).
-This definition of the convolution operation allows to use the existing efficient implementations of 2D convolutions, saving both memory and runtime.
-Lorentz-Core Bottleneck Block
-To build on the concept of hybrid hyperbolic encoders introduced by (Bdeir et al., 2024), we developed the Lorentz Core Bottleneck blocks for hyperbolic ResNet-based models.
-These blocks are similar to standard Euclidean bottleneck blocks but replace the internal 3x3 convolutional layer with our efficient convolutional layer as illustrated in figure 2.
-This design allows us to incorporate hyperbolic structuring of the embeddings within each block while retaining the flexibility and computational efficiency of Euclidean models.
-We interpret this integration as a form of hyperbolic bias, enabling ResNets to leverage hyperbolic representations without requiring fully hyperbolic modeling.
-Euclidean 1x1 ConvManifold Switcher + ScaleHyperbolic 3x3 ConvManifold Switcher + ScaleEuclidean 1x1 ConvBN + ReLUHyper BN + ReLU
-Figure 2: Lorentz-Core Bottleneck Block
-4 Experiments
+where $D(\bm{x},\overline{\bm{0}})_{rescaled}^{k}$ is the scaled distances by plugging in $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Equation 3.
+The time component of the vector is recalculated based on the norm of its spatial components, providing a complete scaling operation for the Lorentz space.
+We apply this tanh scaling during transitions between manifolds, including from Euclidean to Lorentz spaces and between Lorentz spaces of different curvatures.
+It is also used after Lorentz Boosts and direct Lorentz concatenations (Qu & Zou, 2022), as well as following variance-based rescaling in the batch normalization layer, where adjustments can push points outside the valid radius, often resulting in invalid values.
+4 Domain Adaption
+In order to empirically prove the effectiveness of our proposed solutions, we apply them to metric learning, EEG classification, image classification, and image generation tasks.
+In the following, we tackle each of these tasks by offering a brief overview of the problem setting with an existing reference model, our hyperbolic adaptation for this problem, and the goal from this experimental setup.
 4.1 Hierarchical Metric Learning Problem
+Problem Setting and Reference Model
 In their paper Kim et al.
@@
 In the following experiment, we extend HIER to the Lorentz model (LHIER) and compare against the results provided by Kim et al. (2023).
+Moving to Lorentz Space
+To adapt the HIER model to the hyperboloid, we first replace the Euclidean layer norm and linear layer with a Lorentzian norm and linear layer (Bdeir et al., 2024).
+We then modify the HIER loss by setting the hierarchical proxies as learnable hyperbolic parameters and replace the Poincaré distance with the Lorentzain distance.
+This allows us to optimize the hyperbolic parameters directly on the manifold using our new optimization schema.
+We also apply distance rescaling before layer norm and after the linear layer.
+We refer to this setup as LHIER.
+Experimental Goals
+Kim et al.
+(2023) employ the Euclidean AdamW optimizer with Euclidean parameterizations of hyperbolic proxies.
+Their experiments use FP16 precision.
+As the setting is already hyperbolic, significant gains from transitioning to Lorentz space are unlikely.
+We use this setup to evaluate our components’ ability to learn curvature in low-precision, unstable environments.
+Any improvements likely stem from better curvature adaptation to the data and task.
 Experimental Setup
@@
 (2023) and rely on four main datasets: CUB-200-2011 (CUB)(Welinder et al., 2010), Cars-196 (Cars)(Krause et al., 2013), Stanford Online Product (SOP)(Song et al., 2016), and In-shop Clothes Retrieval (InShop)(Liu et al., 2016).
-Performance is measured using Recall@k which is the fraction of queries with one or more relevant samples in their k-nearest neighbors.
-Additionally, all model backbones are pre-trained on Imagenet to ensure fair comparisons with the previous work.
-Moving to Lorentz Space
-To adapt the HIER model to the hyperboloid we first replace the Euclidean linear layer with a Lorentzian linear layer in the model neck and implement our max distance scaling operation after.
-We then set the hierarchical proxies as learnable hyperbolic parameters and optimize them directly on the manifold using our Lorentzian AdamW.
-Finally, we change the Poincaré distance to the Lorentz distance for the LHIER loss and set the hierarchical proxies to be scaled beforehand.
-We continue to use FP16 precision during curvature learning to evaluate the stability of the new optimization scheme.
-Our experiments also include both CNN-based and transformer models to assess whether differences in architecture affect the performance of the hyperbolic heads.
+Performance is measured using Recall@k, the fraction of queries with at least one relevant sample in their k-nearest neighbors.
+All model backbones are pre-trained on ImageNet for fair comparisons.
 Results
-As shown in table 1, our HIER+ model manages to improve the performance of mainly the CNN-based models with improvements ranging in recall@1 from 1-2.3%.
-In contrast, transformer models exhibit marginal gains in most cases and even show slight declines in other scenarios.
-This indicates that the proposed components do not better fully exploit the advantages of transformer architectures compared to other hyperbolic methods.
-The reasons may include architectural limitations or insufficient hyperparameter tuning.
-Notably, Kim et al.
-(2023) employed an extensive set of hyperparameters for model training, which we kept at default values to ensure fair comparisons and avoid costly hyperparameter searches.
 Table 1: Performance of metric learning methods on the four datasets as provided by (Kim et al., 2023).
@@
 Superscripts denote their embedding dimensions and $\dagger$ indicates models using larger input images.
-As in (Kim et al., 2023), network architectures are abbreviated as, R–ResNet50 (He et al., 2016), B–Inception with BatchNorm (Ioffe & Szegedy, 2015), De–DeiT (Touvron et al., 2021), DN–DINO (Caron et al., 2021) and V–ViT (Dosovitskiy et al., 2021)
+As in (Kim et al., 2023), network architectures are abbreviated as, R–ResNet50 (Wang et al., 2017), B–Inception with BatchNorm (Ioffe & Szegedy, 2015), De–DeiT (Touvron et al., 2021), DN–DINO (Caron et al., 2021) and V–ViT (Dosovitskiy et al., 2020)
 Methods Arch. CUB Cars SOP
-R@1 R@2 R@4 R@1 R@2 R@4 R@1 R@10 R@100
-Backbone architecture: CNN
-NSoftmax (Zhai & Wu, 2018) R128 56.5 69.6 79.9 81.6 88.7 93.4 75.2 88.7 95.2
-MIC (Roth et al., 2019) R128 66.1 76.8 85.6 82.6 89.1 93.2 77.2 89.4 94.6
-XBM (Wang et al., 2020) R128 - - - - - - 80.6 91.6 96.2
-XBM (Wang et al., 2020) B512 65.8 75.9 84.0 82.0 88.7 93.1 79.5 90.8 96.1
-HTL (Ge et al., 2018) B512 57.1 68.8 78.7 81.4 88.0 92.7 74.8 88.3 94.8
-MS (Wang et al., 2019) B512 65.7 77.0 86.3 84.1 90.4 94.0 78.2 90.5 96.0
-SoftTriple (Qian et al., 2019) B512 65.4 76.4 84.5 84.5 90.7 94.5 78.6 86.6 91.8
-PA (Kim et al., 2020) B512 68.4 79.2 86.8 86.1 91.7 95.0 79.1 90.8 96.2
-NSoftmax (Zhai & Wu, 2018) R512 61.3 73.9 83.5 84.2 90.4 94.4 78.2 90.6 96.2
-ProxyNCA++ (Teh et al., 2020) R512 69.0 79.8 87.3 86.5 92.5 95.7 80.7 92.0 96.7
-Hyp (Ermolov et al., 2022) R512 65.5 76.2 84.9 81.9 88.8 93.1 79.9 91.5 96.5
-HIER (Kim et al., 2023) R512 70.1 79.4 86.9 88.2 93.0 95.6 80.2 91.5 96.6
-LHIER (ours) R512 72.4 81.5 88.4 89.1 93.5 96.1 81.3 92.1 96.8
-Backbone architecture: ViT
-[…]
-Hyp (Ermolov et al., 2022) De128 74.7 84.5 90.1 82.1 89.1 93.4 83.0 93.4 97.5
-HIER (Kim et al., 2023) De128 75.2 84.2 90.0 85.1 91.1 95.1 82.5 92.7 97.0
-LHIER (ours) De128 75.5 84.7 90.6 85.4 91.7 95.6 82.7 93.5 97.4
-Hyp (Ermolov et al., 2022) DN128 78.3 86.0 91.2 86.0 91.9 95.2 84.6 94.1 97.7
-HIER (Kim et al., 2023) DN128 78.5 86.7 91.5 88.4 93.3 95.9 84.9 94.2 97.5
-LHIER (ours) DN128 78.8 87.0 91.9 88.9 93.2 96.4 85.1 94.9 98.2
-Hyp (Ermolov et al., 2022) V128 84.0 90.2 94.2 82.7 89.7 93.9 85.5 94.9 98.1
-HIER  (Kim et al., 2023) V128 84.2 90.1 93.7 86.4 91.9 95.1 85.6 94.6 97.8
-LHIER (ours) V128 84.6 90.2 93.9 86.7 92.0 95.3 85.9 95.0 98.0
-[…]
-DeiT-S (Touvron et al., 2021) De384 70.6 81.3 88.7 52.8 65.1 76.2 58.3 73.9 85.9
-Hyp (Ermolov et al., 2022) De384 77.8 86.6 91.9 86.4 92.2 95.5 83.3 93.5 97.4
-HIER (Kim et al., 2023) De384 78.7 86.8 92.0 88.9 93.9 96.6 83.0 93.1 97.2
-LHIER (ours) De384 78.3 86.2 91.8 88.7 93.4 96.4 82.8 92.9 96.9
-DINO (Caron et al., 2021) DN384 70.8 81.1 88.8 42.9 53.9 64.2 63.4 78.1 88.3
-Hyp (Ermolov et al., 2022) DN384 80.9 87.6 92.4 89.2 94.1 96.7 85.1 94.4 97.8
-HIER  (Kim et al., 2023) DN384 81.1 88.2 93.3 91.3 95.2 97.1 85.7 94.6 97.8
-LHIER (ours) DN384 81.3 88.4 93.3 91.5 95.1 97.6 85.9 95.0 98.1
-ViT-S (Dosovitskiy et al., 2021) V384 83.1 90.4 94.4 47.8 60.2 72.2 62.1 77.7 89.0
-Hyp (Ermolov et al., 2022) V384 85.6 91.4 94.8 86.5 92.1 95.3 85.9 94.9 98.1
-HIER  (Kim et al., 2023) V384 85.7 91.3 94.4 88.3 93.2 96.1 86.1 95.0 98.0
-LHIER (ours) V384 86.2 92.1 95.2 88.6 93.6 96.3 86.7 95.3 98.3
-4.2 Standard Classification Problem
+R@1 R@2 R@1 R@2 R@1 R@10
+CNN Backbone
+NSoftmax R128 56.5 69.6 81.6 88.7 75.2 88.7
+MIC R128 66.1 76.8 82.6 89.1 77.2 89.4
+XBM R128 - - - - 80.6 91.6
+XBM B512 65.8 75.9 82.0 88.7 79.5 90.8
+HTL B512 57.1 68.8 81.4 88.0 74.8 88.3
+MS B512 65.7 77.0 84.1 90.4 78.2 90.5
+SoftTriple B512 65.4 76.4 84.5 90.7 78.6 86.6
+PA B512 68.4 79.2 86.1 91.7 79.1 90.8
+NSoftmax R512 61.3 73.9 84.2 90.4 78.2 90.6
+ProxyNCA++ R512 69.0 79.8 86.5 92.5 80.7 92.0
+Hyp R512 65.5 76.2 81.9 88.8 79.9 91.5
+HIER R512 70.1 79.4 88.2 93.0 80.2 91.5
+LHIER R512 73.4 82.4 90.0 94.0 81.9 93.1
+Increase % 4.71 3.78 2.04 1.08 1.49 1.20
+As shown in Table 1, LHIER learns curvature without stability issues.
+Our approach improves model performance, with recall@1 gains ranging from 1.49% to 4.71%.
+Additional ablation experiments with different optimizers and configurations are detailed in Table 7.
+4.2 EEG Classification Problem
+Problem Setting
+Given EEG recordings labeled into distinct categories, the goal is to classify new recordings based on signal patterns.
+Each recording is a time series $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the time length.
+Each label is $y\in\{1,\dots,K\}$.
+Given $N$ labeled recordings $(X_{1},y_{1}),\dots,(X_{N},y_{N})$ from an unknown distribution $p$, the task is to train a model $\hat{y}$ that maps EEG signals $X$ to their correct class.
+Reference Model
+The reference model MAtt (Pan et al., 2022) has already taken steps towards improving data preparation.
+In the feature extraction phase, MAtt uses 2 convolutions along the time and sensor dimensions respectively.
+The embeddings are transformed into Symmetric Positive Definite (SPD) data by mapping the feature embeddings from Euclidean space onto the SPD manifold.
+The resulting sequence is then fed into a Manifold Attention Model and the outputs are projected back to the Euclidean space and classified using a fully connected layer.
+Moving to Lorentz Space
+To adapt Matt to a Lorentz space setting, we project the embeddings from the SPD manifold to the Lorentz manifold after the input slicing and sample covariance matrix projection.
+We then pass them through an adapted Lorentz Attention layer that replaces the SPD layers, distance function, and centroid with the Lorentzian counterparts.
+Finally, we use the Lorentz MLR (Bdeir et al., 2024) as a final classification head.
+We refer to this enhanced model as HyperMAtt.
+Experimental Goals
+While this process already represents EEG sequences in Riemannian space, we enhance it further with our Riemannian AdamW Optimizer.
+Hyperbolic spaces are prone to overfitting, especially in EEG classification, where models are trained per subject with limited sessions, leading to data scarcity.
+The proposed optimizer improves regularization, ensuring smoother and more balanced convergence.
+We avoid curvature learning, as it could worsen overfitting and reduce performance, but retain distance rescaling functions.
+Experimental Setup
+We compare our model HyperMAtt on three common EEG datasets for the domain of EEG classification.
+These datasets are MI (Motor Imagery) (Brunner et al., 2008), SSVEP (Steady-State Visual Evoked Potentials) (Nikolopoulos, 2021) and ERN (Error-Related Negativity) (Margaux et al., 2012) representing three different tasks, motor imagery, visual stimuli, and error recognition, respectively.
+Here, we follow the same preprocessing procedure and train/val/test splits as MAtt.
+For the datasets MI and SSVEP, we use Accuracy as an evaluation metric and AUC.
+For the application domain of EEG, we evaluate our method against current state-of-the-art models for EEG Classification including Inception (Burchert et al., 2024), MAtt (Pan et al., 2022), MBEEGSE (Altuwaijri et al., 2022), FBCNet (Mane et al., 2021), TCNet-Fusion (Musallam et al., 2021), EEG-TCNet (Ingolfsson et al., 2020), SCCNet (Wei et al., 2019), EEGNet (Lawhern et al., 2018), and ShallowConvNet (Schirrmeister et al., 2017).
+Results
+In Table 2 we show the performance of HyperMAtt compared to the current state-of-the-art baselines, where the results were aggregated from (Pan et al., 2022; Burchert et al., 2024).
+The performance for HyperMAtt is measured across 3 runs and we report the mean and standard deviation.
+Here, our new optimizer archives state-of-the-art results for SSVEP and ERN, the two comparatively smaller datasets, which are more prone to overfitting.
+We also show the improved performance of our RAdamW optimizer vs the existing RAdam by training the model using both and comparing the results in Table 8.
+We also achieve these results in significantly less time/epoch compared to MAtt.
+Specifically for MI 0.077s/epoch vs 3.4s/epoch, for SSVEP 0.081s/epoch vs 4s/epochs, and for ERN 0.061s/epoch vs 1.9s/epoch.
+Table 2: Performance comparison for the EEG datasets MI, SSVEP, and ERN. We report the average accuracy for MI and SSVEP and the AUC for ERN. The best result is highlighted in bold.
+Models MI SSVEP ERN ShallowConvNet 61.84$\pm$6.39 56.93$\pm$6.97 71.86$\pm$2.64 EEGNet 57.43$\pm$6.25 53.72$\pm$7.23 74.28$\pm$2.47 SCCNet 71.95$\pm$5.05 62.11$\pm$7.70 70.93$\pm$2.31 EEG-TCNet 67.09$\pm$4.66 55.45$\pm$7.66 77.05$\pm$2.46 TCNet-Fusion 56.52$\pm$3.07 45.00$\pm$6.45 70.46$\pm$2.94 FBCNet 71.45$\pm$4.45 53.09$\pm$5.67 60.47$\pm$3.06 MBEEGSE 64.58$\pm$6.07 56.45$\pm$7.27 75.46$\pm$2.34 Inception 62.85$\pm$3.21 62.71$\pm$2.95 73.55$\pm$5.08 MAtt 74.71$\pm$5.01 65.50$\pm$8.20 76.01$\pm$2.28 HyperMAtt (ours) 74.12$\pm$2.91 68.10$\pm$2.41 78.01$\pm$1.30 Increase in % -0.79 3.97 1.25
+4.3 Standard Image Classification Problem
+Problem Setting and Reference Model
+In their work, Bdeir et al.
+(2024) proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al.
+(2022).
+This approach ensured that the convolution outputs remained on the hyperboloid.
+However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions.
+As such, the authors proposed a two versions of their hyperbolic ResNet classifier.
+A fully hyperbolic model with all Lorentz ResNet blocks (HCNN) and a hybrid encoder model with alternating Euclidean and hyperbolic blocks (HECNN).
+Reference Optimization
+To adapt our methods for both models, we introduce the distance rescaling function after every convolution in the encoder.
+Additionally, we attempt to improve the performance by performing a parametrization trick on existing CUDA convolutions to ensure hyperbolic outputs.
+This is done by parametrizing the convolution weight as a rotation operation before passing the input, and then applying a boost operation afterwards.
+The end result is then equivalent to the convolution operation by Bdeir et al.
+(2024) while mitigating computation complexity.
+Specific implementation details can be found in Appendix B.
+We denote our Hybrid and Fully hyperbolic models as HECNN+ and HCNN+ respectively.
+Experimental Goals
+In their paper, (Bdeir et al., 2024) notice lower performance in the fully hyperbolic models compared to the hybrid models and attribute this to instability in training.
+Additionally, they cite crashes and model divergence when learning the curvature instead of setting it to constant.
+This would then be an ideal setting to test the optimization schema and rescaling function and their impact on performance and stability.
+We do not use the RAdamW optimizer in this setting as RSGD is already used and employs proper L2-regularization.
 Experimental Setup
@@
 (2024) and rely on three main datasets: CIFAR10, CIFAR100, and Mini-Imagenet.
-We denote HCNN+ as the fully hyperbolic model from Bdeir et al.
-(2024), updated with efficient convolution implementations and our proposed scaling function.
-This experiment aims to boost model efficiency and address the performance inconsistencies observed in the original version.
 For ResNet-18, we employ HECNN+, following the structure outlined in Bdeir et al.
 (2024) with alternating Euclidean and hyperbolic blocks, and integrate our efficient convolution operation.
-While we do not expect similar efficiency gains to those seen in fully hyperbolic models, our focus here is on enhancing scalability.
-In the case of ResNet-50, we adapt HECNN+ by replacing all blocks with the Lorentz-Core bottleneck block, where we anticipate the most significant efficiency improvements due to the larger model size.
-Our goal is to evaluate if the Lorentz-Core block can maintain strong performance under these conditions.
-For both ResNet-18 and ResNet-50, we use Riemannian SGD with our improved learning scheme and curvature learning.
-We also decouple the encoder and decoder manifolds, allowing each to independently learn its own curvature for enhanced model flexibility.
+We decouple the encoder and decoder manifolds, allowing each to independently learn its own curvature for enhanced model flexibility.
+We also implement ResNet-50 models of the proposed hyperbolic hybrid architectures using similarly constructed bottleneck blocks.
 Results
-For the ResNet-18 experiments, Table 2 shows that the new architectures achieve comparable or better results across almost all cases.
-The smallest improvements are observed with the hybrid models, likely due to the reduced impact of the hyperbolic components compared to a fully hyperbolic model.
-However, we see a significant performance gap between the fully hyperbolic models, where our proposed architecture now matches the performance of the hybrid encoders.
-We hypothesize that the improved scaling helps mitigate the previous performance inconsistencies, with the whole model being two times more memory efficient.
-Table 2: Classification accuracy (%) of ResNet-18 models. We estimate the mean and standard deviation from five runs. The best performance is highlighted in bold (higher is better).
-CIFAR-10 CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.26)$ $(\delta_{rel}=0.23)$ For Cifar100 $(\delta_{rel}=0.20)$ Euclidean (He et al., 2016) ${95.14_{\pm 0.12}}$ $77.72_{\pm 0.15}$ 1.2GB - 12s $65.19_{\pm 0.12}$ Hybrid Poincaré (Guo et al., 2022) $95.04_{\pm 0.13}$ $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ Hybrid Lorentz (Bdeir et al., 2024) ${94.98_{\pm 0.12}}$ ${78.03_{\pm 0.21}}$ ${65.63_{\pm 0.10}}$ Poincaré ResNet (van Spengler et al., 2023) $94.51_{\pm 0.15}$ $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ HECNN Lorentz (Bdeir et al., 2024) ${95.16_{\pm 0.11}}$ ${78.76_{\pm 0.24}}$ 4.3GB - 100s ${65.96_{\pm 0.18}}$ HECNN+ (ours) ${95.15_{\pm 0.07}}$ ${78.80_{\pm 0.12}}$ 3GB - 80s ${65.98_{\pm 0.11}}$ HCNN Lorentz (Bdeir et al., 2024) ${95.15_{\pm 0.08}}$ ${78.07_{\pm 0.17}}$ 10GB - 175s ${65.71_{\pm 0.13}}$ HCNN+ (ours) ${95.17_{\pm 0.09}}$ $\bm{78.81_{\pm 0.19}}$ 5GB - 140s $\bm{66.12_{\pm 0.14}}$
-In the ResNet-50 experiments in Table 3, we observe that HECNN+ significantly outperforms the Euclidean model across all datasets, including Tiny-ImageNet, where other models typically begin to show a drop in accuracy.
-This improvement is likely due to the tighter integration of hyperbolic components and the enhanced scaling, which helps manage the challenges of higher-dimensional embeddings.
-We evaluate the impact of our efficient convolution and Lorentz-Core block in Table 3.
-We see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime.
-We attribute this improvement to the efficient closed-source convolution operations we can now leverage.
-Similar gains are also observed in generation tasks and classification with smaller models.
-However, there remains significant room for further optimization when compared to the Euclidean baseline.
-A key remaining performance bottleneck identified is the batch normalization step, which contributes approximately $60\%$ of the runtime and around $30\%$ of the memory usage.
-The next step would be to factorize the extensive parallel transports and tangent mappings involved in this operation, potentially alleviating the associated overhead.
-Table 3: Classification accuracy (%) of ResNet-50 models. The best performance is highlighted in bold (higher is better).
-CIFAR-10 CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.26)$ $(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ Euclidean (He et al., 2016) 95.1495.14{95.14}95.14 $78.52$ 4.5GB - 30s $66.23$ Hybrid Lorentz (Bdeir et al., 2024) 95.3895.38{95.38}95.38 ${79.35}$ 66.0166.01{66.01}66.01 HECNN (Bdeir et al., 2024) 95.4295.42{95.42}95.42 79.8379.83{79.83}79.83 15.6GB - 300s $66.30$ HCNN+ w Bottleneck Conv.
-(ours) $\bm{95.46}$ $\bm{80.86}$ 8.1GB - 100s $\bm{67.18}$
+For the ResNet-18 experiments, Table 3 shows that the new models are able to remain stable while learning the curvature.
+Additionally, we see significant performance improvements between HCNN and HCNN+, where our proposed architecture now matches the performance of the hybrid encoders.
+We hypothesize that the improved scaling function helps mitigate the previous performance inconsistencies.
+However, the performance difference between the hybrid models is not significant, we hypothesize this is due to the alternating architecture which could limit the effect and instability of hyperbolic components.
+Finally, both models manage to maintain performance while reducing the memory footprint and runtime by approximately $25-50\%$ and $18-25\%$ respectively.
+One other benefit that we find from learning the curvature is quicker convergence where the model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.
+Table 3: Runtime Analysis for the ResNet-18 models.
+We report classification accuracy (%) and estimate the mean and standard deviation from five runs.
+The best performance is highlighted in bold (higher is better).
+CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ For Cifar100 $(\delta_{rel}=0.20)$ Euclidean $77.72_{\pm 0.15}$ 1.2GB - 12s $65.19_{\pm 0.12}$ Hybrid ($\mathbb{P}$) $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ Hybrid ($\mathbb{L}$) ${78.03_{\pm 0.21}}$ ${65.63_{\pm 0.10}}$ ResNet ($\mathbb{P}$) $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ HECNN ($\mathbb{L}$) ${78.76_{\pm 0.24}}$ 4.3GB - 100s ${65.96_{\pm 0.18}}$ HECNN+ ${78.80_{\pm 0.12}}$ 3GB - 80s ${65.98_{\pm 0.11}}$ HCNN ($\mathbb{L}$) ${78.07_{\pm 0.17}}$ 10GB - 175s ${65.71_{\pm 0.13}}$ HCNN+ $\bm{78.81_{\pm 0.19}}$ 5GB - 140s $\bm{66.12_{\pm 0.14}}$
+In the ResNet-50 experiments in Table 4, HECNN+ significantly outperforms both the Euclidean model and the base hybrid model across both datasets.
+We also see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime.
+We attribute this improvement to efficient closed-source convolution operations we can now leverage.
+Table 4: Runtime Analysis for ResNet-50 models. We report classification accuracy (%) and the best performance is highlighted in bold (higher is better).
+CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ Euclid $78.52$ 4.5GB - 30s $66.23$ HECNN 79.8379.83{79.83}79.83 15.6GB - 300s $66.30$ HECNN+ $\bm{80.86}$ 8.1GB - 100s $\bm{67.18}$
 4.4 VAE Image Generation
-Experimental Setup
+Problem Setting and Reference Model
 Previous research has highlighted the ability of hyperbolic neural networks to perform better when using lower dimensional embeddings due to the more expressive space (Nagano et al., 2019; Mathieu et al., 2019; Ovinnikov, 2019; Hsu et al., 2020).
@@
 (2024) perform this study on image generation and reconstruction for their fully hyperbolic models.
-In the following section, we reproduce the experimental setup and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers.
+Experimental Setup
+We reproduce the experimental setup from (Bdeir et al., 2024) and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers.
 We also use curvature learning with our adjusted Riemannian SGD learning scheme.
 Results
-We see in table 4 that our implementation of the fully hyperbolic VAE achieves better performance on both datasets.
-It should also be noted that this is achieved with around 2.5x less memory and at 3x greater training speed.
-This helps demonstrate both the effectiveness of our proposed curvature learning process and our efficient model components.
+Our fully hyperbolic VAE implementation outperforms on both datasets, as shown in Table 5, while using 2.5x less memory and training 3x faster.
+This highlights the effectiveness of our curvature learning process and efficient model components.
 Table 5: Reconstruction and generation FID of manifold VAEs across five runs (lower is better).
-CIFAR-100 CelebA Rec.
-FID Gen.
-FID Rec.
-FID Gen.
-FID Euclidean $63.81_{\pm 0.47}$ $103.54_{\pm 0.84}$ $54.80_{\pm 0.29}$ $79.25_{\pm 0.89}$ Hybrid Poincaré (Mathieu et al., 2019) $62.64_{\pm 0.43}$ $98.19_{\pm 0.57}$ $54.62_{\pm 0.61}$ $81.30_{\pm 0.56}$ Hybrid Lorentz (Nagano et al., 2019) $62.14_{\pm 0.35}$ $98.34_{\pm 0.62}$ $54.64_{\pm 0.34}$ $82.78_{\pm 0.93}$ HCNN Lorentz (Bdeir et al., 2024) $61.44_{\pm 0.64}$ $100.27_{\pm 0.84}$ $54.17_{\pm 0.66}$ $78.11_{\pm 0.95}$ HCNN+ Lorentz (Ours) $57.69_{\pm 0.52}$ $98.14_{\pm 0.44}$ $52.73_{\pm 0.27}$ $77.98_{\pm 0.32}$
+CIFAR-100 CelebA
+Rec. FID Gen. FID Rec. FID Gen. FID
+Euclid $63.81_{\pm 0.47}$ $103.54_{\pm 0.84}$ $54.80_{\pm 0.29}$ $79.25_{\pm 0.89}$
+Hybrid ($\mathbb{P}$) $62.64_{\pm 0.43}$ $98.19_{\pm 0.57}$ $54.62_{\pm 0.61}$ $81.30_{\pm 0.56}$
+Hybrid ($\mathbb{L}$) $62.14_{\pm 0.35}$ $98.34_{\pm 0.62}$ $54.64_{\pm 0.34}$ $82.78_{\pm 0.93}$
+HCNN ($\mathbb{L}$) $\underline{61.44}_{\pm 0.64}$ $\underline{100.27}_{\pm 0.84}$ $\underline{54.17}_{\pm 0.66}$ $\underline{78.11}_{\pm 0.95}$
+HCNN+ ($\mathbb{L}$) $57.69_{\pm 0.52}$ $98.14_{\pm 0.44}$ $52.73_{\pm 0.27}$ $77.98_{\pm 0.32}$
+Decrease in % 6.10 2.12 2.66 0.17
 5 Conclusion
-In our work, we present many new components and schemas for the use of hyperbolic deep learning in hyperbolic vision.
-We test these components in three vision tasks and prove the potential of these new components even in float16 conditions.
-However, there is still significant room for improvement.
-Further optimizations to the batch normalization layers could enhance the efficiency of hyperbolic models.
-Additionally, a key challenge remains with the hyperbolic linear layers when reducing dimensionality.
-Currently, we match norms to simulate a rotation operation, but we encourage exploring alternative approaches that align more naturally with the mathematical properties of the manifold.
-Appendix A Appendix
-A.1 Operations in hyperbolic geometry
+In this work, we propose a robust curvature-aware optimization framework for hyperbolic deep learning, addressing challenges in stability, computational efficiency, and overfitting.
+By introducing Riemannian AdamW, a novel distance rescaling function, and leveraging efficient CUDA implementations, our approach achieved performance improvements in hierarchical metric learning, EEG classification, and image classification tasks, while reducing computational costs, highlighting the practicality of our methods for large-scale applications.
+These findings emphasize the importance of proper curvature adaptation in hyperbolic learning, paving the way for future research in optimizing hyperbolic embeddings across diverse fields.
+6 Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning.
+We are committed to contributing positively to society and human well-being, while respecting the privacy of the subjects involved in our evaluation.
+Our experiments utilize three established EEG datasets: MI, SSVEP, and ERN, which involve data from human subjects.
+These datasets are fully anonymized and do not contain any personally identifiable information.
+Additionally, they are publicly available and widely used within the machine learning community.
+Appendix A Operations in hyperbolic geometry
 Parallel Transport
@@
 $\bm{O}^{+}(1,n)={\mathrm{\textbf{A}}\in\mathbb{R}^{(n+1)\times(n+1)}|\forall\bm{x}\in\mathbb{L}^{n}_{K}:\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{x}\rangle_{\mathbb{L}}=-\frac{1}{K},(\mathrm{\textbf{A}}\bm{x})_{0}>0)}.$ (6)
-Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ Moretti (2002).
+Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ (Moretti, 2002).
 The rotation matrix R is designed to rotate points around the time axis and is defined as
@@
 with $\gamma=\frac{1}{\sqrt{1-||\bm{v}||^{2}}}$. However, this can also be any operation that scales the norms of the space values without changing the vector orientation.
-A.2 Scaling Lorentzian Vectors
+Appendix B Convolution Trick
+To address the computational requirements issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al.
+(2021), which decomposes the transformation into a Lorentz boost and a Lorentz rotation.
+Using this definition, we replace the matrix multiplication employed by Bdeir et al.
+(2024) for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost.
+Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
+To apply this concept to the convolution operation, the convolution weights, after unfolding, must form a rotation matrix.
+We define the dimensions of this matrix as $n=(channels_{in}\cdot kernel_{width}\cdot kernel_{height})$ and $n^{\prime}=channels_{out}$ respectively.
+We then use either rotation operation presented above to a norm-preserving transformation $\bm{z}=\bm{W^{T}}\bm{x}\cdot\frac{\lVert\bm{x}\rVert}{\lVert\bm{W^{T}}\bm{x}\rVert}$ where $\bm{W}\in\mathbb{R}^{(n^{\prime}\cdot n)}$.
+This formulation allows us to utilize existing efficient implementations of the convolution operation by directly parameterizing the kernel weights before passing them into the convolutional layer.
+Finally, we formalize the new Lorentz Convolution as:
+$\bm{out}=\text{LorentzBoost(DistanceRescaling(RotationConvolution(}\bm{x}\textbf{)))}$)) (9)
+where TanhRescaling is the operation described in Eq.4 and RotationConvolution is a normal convolution parameterized through the procedure in Algorithm 2 where Transform is the norm-preserving transformation above.
+Algorithm 2 Lorentz Convolution Parametarization
+  $\bm{W}\in\mathbb{R}^{C_{in},C_{out},K_{width},K_{length}}$
+  function Adaptweight
+     if $K_{width}\cdot K_{length}\cdot C_{in}>C_{out}$ then
+        return $\bm{W}$
+     else
+        $\bm{W}=\bm{W}.reshape(K_{width}\cdot K_{length}\cdot C_{in},C_{out})$
+        $\bm{\hat{W}}_{core}=$ Transform ($\bm{W}$)
+        return $\bm{\hat{W}}.reshape(C_{in},C_{out},K_{width},K_{length})$
+     end if
+  end function
+Appendix C Scaling Lorentzian Vectors
 Tanh Scaling
@@
 $\bm{x}_{s_{rescaled}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}$ (13)
-A.3 Ablations
+Appendix D Ablations
+D.1 Image Classification
 We test the effect of individual model components in Table 6.
@@
 In the case of attempting to learn the curvature without the proposed optimizer schema, the model breaks completely down due to excessive numerical inaccuracies.
-One other benefit that we find from learning the curvature is quicker convergence.
-The model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.
 Table 6: Ablation experiments for Resnet-50 models on Cifar-100.
 CIFAR-100 HCNN+ - Default $\bm{80.86}$ HCNN+ - fixed curve $79.6$ HCNN+ - no scaling $80.13$ HCNN+ - no optim scheme $NaaN$
+D.2 Metric Learning
+In Table 7 we present the LHIER model results under different optimizers and components.
+Table 7: Ablation on components for Metric Learning
+Methods Arch. CUB Cars SOP
+R@1 R@2 R@1 R@2 R@1 R@10
+CNN Backbone
+LHIER RSGD R512 64.2 72.1 73.9 81.4 67.8 73.9
+LHIER RAdam R512 70.1 79.8 87.6 92.0 79.8 90.3
+LHIER CST Curvature R512 71.2 80.8 88.7 91.8 79.1 89.7
+LHIER No Learning Schema R512 65.8 72.0 - - - -
+LHIER R512 73.4 82.4 90.0 94.0 81.9 93.1
+D.3 EEG Classification
+We also train the model with the original Adam Optimizer and include the results in Table 8. This clearly shows the improvement provided by our RAdamW optimizer over standard RAdam.
+Table 8: Performance comparison for the EEG datasets MI, SSVEP, and ERN. We report the average accuracy for MI and SSVEP and the AUC for ERN. The best result is highlighted in bold.
+Models MI SSVEP ERN HyperMAtt + Adam (ours) 68,53$\pm$1.29 62.42$\pm$2.62 74.98$\pm$5.84 HyperMAtt + AdamW (ours) 74.12$\pm$2.91 68.10$\pm$2.41 78.01$\pm$1.30

### v4

@@
 For improved stability, we introduce a novel fine-tunable hyperbolic scaling approach to constrain hyperbolic embeddings and reduce approximation errors.
-Using this along with our curvature-aware learning schema for Lorentzian Optimizers enables the combination of curvature and non-trivialized hyperbolic parameter learning.
-Our approach demonstrates consistent performance improvements across Computer Vision, EEG classification, and hierarchical metric learning tasks achieving state-of-the-art results in two domains and drastically reducing runtime.
-Machine Learning, ICML
+Using this along with our curvature-aware learning schema for Riemannian Optimizers enables the combination of curvature and non-trivialized hyperbolic parameter learning.
+Our approach demonstrates consistent performance improvements across Computer Vision, EEG classification, and hierarchical metric learning tasks while greatly reducing runtime.
+Our code is publicly available at https://github.com/inboxedshoe/Robust-Hyperbolic-Learning
 1 Introduction
@@
 Unlike Euclidean space, hyperbolic space has negative curvature, allowing it to represent exponentially growing distances.
-This makes these manifolds ideal for tasks like natural language processing, graph representation, and recommendation systems (Peng et al., 2022).
-By embedding data into hyperbolic space, models can capture complex relationships more effectively, often with fewer parameters (Ganea et al., 2018), leading to improved performance and interpretability.
-Hyperbolic geometry has also found applications in computer vision, where its ability to better separate embeddings in high-dimensional spaces has shown promise in improving tasks like image classification and segmentation(Bdeir et al., 2024; van Spengler et al., 2023; Khrulkov et al., 2020; Atigh et al., 2022), few-shot learning(Liu et al., 2020), and feature representation (Yang et al., 2023).
-Two main derivations of hyperbolic space have emerged as learning frameworks: the hyperboloid or the Lorentz space ($\mathbb{L}$) and the Poincaré manifold ($\mathbb{P}$).
-Typically, the hyperboloid, or Lorentz space, offers better operational stability, as demonstrated by Mishne et al.
-(2022) but lacks clear definitions for basic vector operations such as addition and subtraction.
-To bridge this gap, recent research has focused on defining Lorentzian versions of common deep learning operations, such as the feed-forward layer (Chen et al., 2022; Dai et al., 2021; Ganea et al., 2018), convolutional layer (Chen et al., 2022; Qu & Zou, 2023; Dai et al., 2021), and MLRs (Bdeir et al., 2024).
+This makes these manifolds ideal for tasks like natural language processing, graph representation, and metric learning [33].
+By embedding data into hyperbolic space, models can capture complex relationships more effectively, and often with fewer parameters [11].
+Hyperbolic geometry has also been applied in computer vision, where its ability to better separate embeddings in high-dimensional spaces has shown promise in improving tasks like image classification and segmentation [3, 40, 18, 2], few-shot learning [24], and feature representation [44].
+These works rely on two main derivations of the hyperbolic space, the hyperboloid or Lorentz space ($\mathbb{L}$) and the Poincaré manifold ($\mathbb{P}$).
+Typically, the hyperboloid offers better operational stability, as demonstrated by Mishne et al.
+[29] but lacks clear definitions for basic vector operations such as addition and subtraction.
+To bridge this gap, recent research has focused on defining Lorentzian variant of common deep learning operations, such as the feed-forward layer [6, 8, 11], the convolutional layer [6, 34, 8], and multinomial linear regression (MLR) [3].
 However, the use of these hyperbolic operations comes with challenges.
-Computations in hyperbolic space are more complex and expensive due to the increased mathematical complexity and the lack of optimized CUDA implementations, which can slow down training and drastically increase memory requirements.
-Additionally, optimization on hyperbolic manifolds can be unstable, especially in low floating-point precision environments.
-This has required research to rely on parameter clamping techniques, which can cause non-smooth gradient updates, to remain within the accurate representation radius for manifolds Mishne et al.
-(2022); Guo et al.
-(2022).
-Otherwise, performance either suffers or training becomes impossible due to the precision errors causing mathematically undefined operations.
-The instability is only exacerbated when we incorporate the negative curvature as a learned parameter since it directly affects the representational radius and requires the projection of vectors between these different curvatured manifolds.
-Finally, in data-scarce scenarios, the higher representational capacity makes the hyperbolic spaces more prone to overfitting which heavily impacts their generalization ability.
+Computations in hyperbolic space are more complex and expensive, and the lack of optimized CUDA implementations drastically slows down training and increases memory requirements.
+Additionally, optimizing hyperbolic parameters such as class prototypes or batchnorm means can be unstable, especially in low precision floating-point environments.
+This has required research to rely on parameter clamping techniques, which can cause non-smooth gradient updates, to remain within the accurate representation radius of the manifolds [29, 16].
+The instability is only exacerbated when we incorporate the negative curvature as a learned parameter since it directly affects the embedding space.
+Finally, in data-scarce scenarios, the higher representational capacity makes the hyperbolic spaces more prone to overfitting, which heavily impacts their generalization ability.
 This work addresses key challenges in the Lorentz model.
 To combat overfitting, we introduce an AdamW-based optimizer that enhances regularization and generalization in hyperbolic learning.
-For learning instability, we propose an optimization framework that stabilizes curvature parameter learning.
+For learning instability, we propose an optimization framework that stabilizes the learning of curvature parameters.
 We also present a smoother scaling function to replace weight clipping, ensuring hyperbolic vectors remain within the representational radius.
-To tackle computational complexity, we introduce an implementation trick that leverages efficient CUDA-based convolutions for hyperbolic learning.
-Improved stability not only boosts model performance but also enables lower-precision training, enhancing efficiency.
-Our contributions are four-fold:
+To address computational complexity, we introduce an implementation trick that leverages efficient CUDA-based convolutions for hyperbolic learning.
+Improved stability not only boosts model performance but also enables lower-precision training, further enhancing efficiency.
+Our contributions are then four-fold:
   1. 1.
-    We propose a formulation for Riemannian AdamW and an alternative schema for Riemannian optimizers that stabilize curvature learning for hyperbolic parameters.
+    We propose an alternative schema for Riemannian optimizers that stabilize curvature learning for hyperbolic parameters and a formulation for the Riemannian AdamW
   2. 2.
@@
   3. 3.
-    We present HCNN+, HyperMAtt, and LHEIR, an application of our proposed optimization scheme for the domains of classical image classification, EEG classification, and hierarchical metric learning respectively.
+    We present LHEIR, HyperMAtt, and HCNN+ as applications of our proposed optimization scheme for the domains of hierarchical metric learning, EEG classification, and computer vision for image classification and generation, respectively.
   4. 4.
-    We empirically show the effectiveness of our proposed methods in three domains, hierarchical metric learning, Image, and EEG classification, where we achieve state-of-the-art performance in two domains with a significant computational speed-up.
-Our code is publicly available at [redacted for double-blind submission].
+    We empirically show the effectiveness of our proposed methods in five domains, hierarchical metric learning, EEG classification, graph embedding, image classification, and image generation to show the effectiveness of our optimizer in different problem settings.
+We improve performance in all domains with a significant computational speed-up.
 2 Related Work
 Hyperbolic Embeddings in Deep Learning
-With the success of employing hyperbolic manifolds in NLP models (Zhu et al., 2021; Dhingra et al., 2018; Tifrea et al., 2018) hyperbolic embeddings have extended to the computer vision domain.
-Initially, many of the works relied on a hybrid architecture, utilizing Euclidean encoders and hyperbolic decoders (Mettes et al., 2023).
-This was mainly due to the high computational cost of hyperbolic operations in the encoder, as well as the lack of well-defined alternatives for Euclidean operations.
-However, this trend has begun to shift towards the utilization of fully hyperbolic encoders as can be seen in the hyperbolic Resnets by Bdeir et al.
-(2024) and van Spengler et al.
-(2023).
-Both works offer hyperbolic definitions for 2D convolutional layer, batch normalization layer, and an MLR for the final classification head.
+Initially, many hyperbolic deep learning methods relied on a hybrid model architecture that utilizes Euclidean encoders and hyperbolic decoders [28].
+Euclidean encoders avoid the high computational complexity of hyperbolic operations, as well as the lack of well-defined hyperbolic alternatives for Euclidean components.
+However, this trend has begun to shift towards fully hyperbolic models.
+Chen et al.
+[6] propose hyperbolic components for a fully connected linear layer, a graph convolution layer, and an attention layer with the square Lorentzian distance as a similarity metric.
+For the classification head, they learn class prototypes directly on the hyperbolic manifold and use the same distance metric as a loss estimator, following similar work in the literature [2, 19, 28].
+This work was further extended to computer vision by Bdeir et al.
+[3] and van Spengler et al.
+[40] with both developing fully hyperbolic ResNets, using Riemannian batch normalization layers that rely on a learnable mean also embedded and optimized in hyperbolic space.
 Bdeir et al.
-(2024) even attempt to hybridize the encoder by employing the Lorentz manifold in blocks that exhibit higher output hyperbolicity.
-While this has led to notable performance improvements, both models suffer from upscaling issues.
-Attempting to replicate these approaches for larger datasets or bigger architectures becomes much less feasible in terms of time and memory requirements.
-Instead, our approach places higher focus on efficient components to leverage the beneficial hyperbolic properties of the model while minimizing the memory and computational footprint.
+[3] attempt to alleviate these issues by including a hybrid encoder that only applies the hyperbolic components in blocks that exhibit higher embedding hyperbolicity.
+Although this has led to notable performance improvements, both models suffer from upscaling issues.
+Attempting to apply these approaches to larger datasets or larger architectures becomes less feasible in terms of time and memory requirements.
+Instead, our approach places a greater focus on efficient components to leverage the beneficial hyperbolic properties of the model while minimizing the memory and computational footprint.
 Curvature Learning
@@
 In their studies, Gu et al.
-(2018) and Giovanni et al.
-(2022) achieve this by using a radial parametrization that implicitly models variable curvature embeddings under an explicitly defined, fixed 1-curve manifold.
-This method enables them to simulate K-curve hyperbolic and spherical operations under constant curvature for the mixed-curve manifold specifically, a combination of the Euclidean, spherical, and Poincaré manifold.
+[14] and Giovanni et al.
+[13] achieve this by using a parametrization that implicitly models variable-curvature embeddings under an explicitly defined 1-curve manifold.
+This method enables them to simulate K-curve hyperbolic and spherical operations under constant curvature.
+They apply this method on mixed-curve manifold embeddings where every portion of the embedding belongs to either the Euclidean, spherical, or Poincaré manifold.
 Other approaches, such as the one by Kochurov et al.
-(2020), simply set the curvature to a learnable parameter but do not account for the manifold changes in the Riemannian optimizers.
-This leads to hyperbolic vectors being updated with mismatched curvatures and others being inaccurately reprojected, resulting in instability and accuracy degradation.
+[20], simply set the curvature to a learnable parameter but do not account for the manifold changes while performing the optimization steps with the Riemannian optimizers.
+This leads to mathematical inconsistencies when updating the hyperbolic parameters, resulting in instability and accuracy degradation.
 Additionally, some methods, like the one by Kim et al.
-(2023), store all manifold parameters as Euclidean vectors and project them before use.
-While this approach partially mitigates the issue of mismatched curvature operations, it remains less accurate and more computationally expensive.
-In comparison, our proposed optimization schema maintains the parameters on the manifold and optimizes them directly by performing the necessary operations to transition between the variable curvature spaces.
+[19], store all manifold parameters as Euclidean vectors and project them before use.
+While this approach partially mitigates the issue of mismatched curvature operations, it requires repeated backpropagation through hyperbolic mappings, making it computationally expensive and more susceptible to projection errors.
+Other methods [3, 40, 6] use a combination of Euclidean parametrization, and direct hyperbolic parameter learning depending on the component and required precision.
 3 Methodology
@@
 There are many conformal models of hyperbolic space but we focus our work on the hyperboloid, or Lorentz manifold.
-The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}=[x_{t},\bm{x_{s}}]\in\mathbb{R}^{n+1}~{}|~{}\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=\frac{-1}{K},~{}x_{t}>0\}$ where $\frac{-1}{K}=c$ and the Lorentzian inner product as Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}.$ It then follows that ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ which is an important characteristic to note for the stability issues presented later on.
-The formulation $\mathcal{L}^{n}$ then presents the Lorentz manifold as the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}=[\sqrt{K},0,\cdots,0]^{T}$.
-We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector, as the space dimension $\bm{x_{s}}$.
-Below are the basic operations presented by the manifold with more complex operations provided in Appendix A.
+The n-dimensional Lorentz model $\mathbb{L}^{n}_{K}=(\mathcal{L}^{n},\mathfrak{g}_{\bm{x}}^{K})$ is defined with $\mathcal{L}^{n}:=\{\bm{x}=[x_{t},\bm{x_{s}}]\in\mathbb{R}^{n+1}~|~\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=-K,~x_{t}>0\}$} implying a negative curvature of $\frac{-1}{K}$ and the Lorentzian inner product as Riemannian metric $\mathfrak{g}_{\bm{x}}^{K}=\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}:=-x_{t}y_{t}+\bm{x}_{s}^{T}\bm{y}_{s}.$ It then follows that $\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}\leq-K$ which is an important characteristic to note for the stability issues presented later on.
+The formulation $\mathcal{L}^{n}$ then presents the Lorentz manifold as the upper sheet of a two-sheeted hyperboloid centered at $\overline{\bm{0}}^{K}=[\sqrt{K},0,\cdots,0]^{T}$.
+We inherit the terminology of special relativity and refer to the first dimension of a Lorentzian vector as the time component $x_{t}$ and the remainder of the vector as the space dimension $\bm{x_{s}}$.
+Below are the basic operations presented by the manifold with more complex operations provided in Appendix˜A in the supplementary material.
 Distance
-Distance in hyperbolic space is the magnitude of the geodesic forming the shortest path between two points. Let $\bm{x},\bm{y}\in\mathbb{L}^{n}_{K}$, the distance between them is given by
-$d_{\mathbb{L}}(\bm{x},\bm{y})=\sqrt{K}\cosh^{-1}(\frac{-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}{K})$ (1)
-We also define the square distance by Law et al. (2019) as $d^{2}_{\mathbb{L}}(\bm{x},\bm{y})=||\bm{x}-\bm{y}||_{\mathbb{L}}^{2}=-2K-2\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$.
+[…]
 Exponential and Logarithmic Maps
-Since the Lorentz space is a Riemannian manifold, it is locally Euclidean.
-This can best be described through the tangent space $\mathbb{T}_{\bm{x}}\mathcal{M}$, a first-order approximation of the manifold at a given point $\bm{x}$.
-The exponential map, $\exp^{K}_{\bm{x}}\left(\bm{z}\right):\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}\rightarrow\mathbb{L}^{n}_{K}$ is then the operation that maps a tangent vector $\mathbb{T}_{\bm{x}}\mathbb{L}_{K}^{n}$ onto the manifold through $\exp_{\bm{x}}^{K}(\bm{z})=\cosh(\alpha)\bm{x}+\sinh(\alpha)\frac{\bm{z}}{\alpha},\text{ with }\alpha=\sqrt{1/K}||\bm{z}||_{\mathbb{L}},~{}~{}||\bm{z}||_{\mathbb{L}}=\sqrt{{\langle\bm{z},\bm{z}\rangle}_{\mathbb{L}}}.$ The logarithmic map is the inverse of this mapping and can be described as $\log_{\bm{x}}^{K}(\bm{y})=\frac{\cosh^{-1}(\beta)}{\sqrt{\beta^{2}-1}}\cdot(\bm{y}-\beta\bm{x}),\text{ with }\beta=-\frac{1}{K}\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}.$
+[…]
 3.2 Riemannian Optimization Schema
 Background
-In the GeoOpt hyperbolic learning library, Kochurov et al.
-(2020) introduces the curvature $c$ of hyperbolic space as a learnable parameter.
-However, no subsequent work has fully taken advantage of this feature despite the authors presenting ready implementations of hyperbolic operations that account for this curvature.
-Our empirical testing also reveals that naively learning this parameter often leads to instability and performance decline.
-We attribute these issues to the simplistic curvature updates in Riemannian optimizers, which neglect to adjust hyperbolic operations, weights, and gradients after updating the curvature.
-In the current Riemannian optimization approach, Euclidean gradients and momentums are projected onto the manifold to update hyperbolic parameters.
-However, since optimization depends on curvature, adjusting it during learning makes previous projections, gradients, and parameters inconsistent with the new geometry.
-Essentially, the current schema applies updates using the new curvature on vectors defined by the old curvature, leading to mismatches.
-To illustrate this, LABEL:fig:curve_inconsistencies shows two manifolds with different curvatures.
-It is evident that points defined on the first manifold do not lie on the second, and the tangent planes of the first manifold are not aligned with those of the second.
-Thus, using the new curvature operations on points from the old model is not mathematically sound practice.
-This inconsistency introduces instability, undermining the reliability of curvature learning and often leading to undefined values during the parameter updates.
-One example of this is the distance function Equation 1 which requires the bound ${}\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}<\frac{-1}{K}$ to hold.
-If the new curvature parametrization at time t $K_{t}>K_{t-1}$ then the input into the inverse hyperbolic cosine has a lower bound less than 1, leaving room for the operation to be mathematically undefined.
-While GeoOpt does not address this particular optimization challenge, it does implement an ”N-stabilize step”.
-This step aims to reduce instability by recalculating the time component of the hyperbolic parameters every N-steps to ensure they remain on the manifold.
-However, this approach is done after the update step and does not prevent invalid parameter values during optimization.
+Many existing hyperbolic models adopt hybrid learning approaches that combine Euclidean parameterization with direct hyperbolic optimization.
+Our work focuses on the latter.
+This motivates our reliance on GeoOpt [20], a widely adopted library for Riemannian optimization based on the definitions by Bécigneul and Ganea [4], whose methodology is also reflected in related work.
+Here, the curvature $K$ of the hyperbolic space is set as a learnable parameter.
+However, empirically, we find that naively optimizing this curvature often leads to instability and performance degradation.
+We argue that this stems from a fundamental oversight: Riemannian optimizers update curvature without adjusting dependent hyperbolic operations, weights, or gradients, creating geometric inconsistencies.
+Given a hybrid model with both Euclidean and hyperbolic parameters, we define the set of learnable parameters as $\bm{\theta}=[\bm{\theta}_{\mathbb{E}},\bm{\theta}^{K}_{\mathbb{L}},\bm{\theta}_{K}]$, where $\bm{\theta}_{\mathbb{E}}$ are the parameters optimized in Euclidean space, $\bm{\theta}^{K}_{\mathbb{L}}$ are the parameters constrained to a Riemannian manifold $\mathcal{M}$ with curvature $K$ and origin $\overline{\bm{0}}^{K}$, and $\bm{\theta}_{K}$ are learnable curvature parameters for the manifold.
+When optimizing the curvature, we refer to $\bm{\theta}^{K_{t}}_{\mathbb{L}}$ as the hyperbolic parameters defined on the manifold with the value of the curvature parameter at timestep $t$.
+It is easy to see here that $\bm{\theta}^{K}_{\mathbb{L}}$ is dependent on $\bm{\theta}_{K}$ since the hyperbolic parameters are defined by their values.
+Current Riemannian optimization first calculates the Euclidean gradient $\mathcal{G}$ through normal backpropagation.
+If the parameter we are currently updating is Euclidean ($\bm{\theta}_{\mathbb{E}}$), we perform a typical Euclidean update step.
+If the parameter being updated is hyperbolic ($\bm{\theta}^{K}_{\mathbb{L}}$), the optimizer projects the gradient onto the tangent space of the parameter $\mathcal{T}_{\bm{\theta}_{\mathbb{L}}}$ in the process egrad2rgrad described in Appendix A.
+Momentum vectors typically used with optimizers like Adam and SGD are also initialized and updated on the tangent space.
+Finally, the update step is performed using some form of retraction or exponential map.
+Currently, Riemannian optimizers treat $\bm{\theta}_{K}$ as a Euclidean parameter.
+However, since the curvature defines the geometry of the manifold, changing it during training renders prior projections, gradient momentums, and $\bm{\theta}^{K}_{\mathbb{L}}$ parameters misaligned with the new curvature.
+GeoOpt attempts to mitigate this instability through an "N-stabilize step," which periodically recomputes the time components of hyperbolic parameters to ensure adherence to the manifold.
+However, this occurs after updates, failing to prevent invalid intermediate states during optimization.
+As a concrete example, let $H$ and $H^{\prime}$ be two Lorentz manifolds with curvatures $K$ and $K^{\prime}$.
+Let $\mathbf{x}$ be a learnable model parameter that lies on $H$, meaning it satisfies $\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=-K$.
+The hyperbolic distance from this point to itself, $d_{\mathcal{L}}(\mathbf{x},\mathbf{x})$, should always be 0.
+However, if the optimizer first updates the curvature to $K^{\prime}$ but does not yet update the parameter $\mathbf{x}$, any subsequent operation will use the new curvature $K^{\prime}$ with the old parameter coordinates.
+The distance calculation becomes:
+$d_{\mathcal{L}}(\mathbf{x},\mathbf{x})=\sqrt{K^{\prime}}\cdot\operatorname{acosh}\left(\frac{-\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}}{K^{\prime}}\right)=\sqrt{K^{\prime}}\cdot\operatorname{acosh}\left(\frac{K}{K^{\prime}}\right)$
+If $K<K^{\prime}$, the term $\frac{K}{K^{\prime}}$ is less than 1.
+The $\operatorname{acosh}$ function is undefined for inputs less than 1, leading to NaN values and a model crash.
+If $K>K^{\prime}$, the calculation does not crash, but the distance is no longer 0.
+This incorrect distance creates invalid gradients, which destabilizes training.
+Previous similar works relied on rigorous training regimes with separate optimizers for the curvature, very small learning rates, and "burn-in" epochs [36, 7] which we attempt to avoid.
+Other solutions such as clipping may prevent undefined operations but cannot easily account for the inconsistencies in the second scenario.
 Curvature Aware Optimization
-To address this, we propose a method that tracks both the old and new curvature parameter values $\bm{\theta}_{K}$ throughout updates.
-Specifically, we propose a simple approach where all updates are first applied under the old curvature, then the new curvature is derived and finally, the hyperbolic vectors are mapped from the old manifold to the new one.
-The previously proposed mapping technique in the N-stabilize step alters the relative magnitudes of the hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ and gradients $\mathcal{G}$, as well as the directions of the momentums, which can degrade performance and lead to training instability.
-Instead, we recommend projecting points onto the tangent space at the origin of the old curvature $\overline{\bm{0}}_{t-1}$ using the logarithmic map and then projecting back after the curvature update.
-Gradients and their momentum can be similarly projected or parallel transported Appendix A) to the tangent space.
-The entire optimization schema can then be seen in Algorithm 1.
-Here, we choose to apply parallel transport, or any similar approximation of it, because it preserves the gradient vectors’ directions which is important for the optimization procedure.
-We also choose to use the tangent space at the origin as the mapping intermediate since it remains relatively stable when the curvature is updated, as shown in LABEL:fig:curve_inconsistencies.
-Finally, we prove that the vector norms and their hyperbolic distances to the origin are preserved since $D(\exp^{K_{1}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{1}}=||y||=D(\exp^{K_{2}}_{\overline{\bm{0}}}\left(\bm{y}\right),\overline{\bm{0}})^{K_{2}}$ where $\bm{y}\in\mathcal{T}_{\overline{\bm{0}}}\mathcal{M}$.
+To address this, we propose a method that groups parameters based on their respective spaces and staggers their updates.
+Specifically, at timestep $t$ we first isolate $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}$ and update these parameters first.
+We then update $\bm{\theta}_{\mathbb{E}}$ followed by $\bm{\theta}_{K}$.
+At this stage, we have updated all the model parameters, but $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}$ are still defined on the old curvature, so we must use a mapping function to update $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}\rightarrow\bm{\theta}^{K_{t}}_{\mathbb{L}}$.
+The N-stabilize step used by GeoOpt can be seen as a pseudo-map, but it alters the relative magnitudes of the hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ and gradients $\mathcal{G}$, as well as the directions of the momentums, which can degrade performance and lead to training instability.
+As such, we identify two alternate mapping techniques commonly used in the literature and compare their properties.
+We start with the scaling function used by Tabaghi et al.
+[38], Skopek et al.
+[36].
+In their works the authors simply multiply $\bm{\theta}^{K_{t-1}}_{\mathbb{L}}$ by the scaling factor $l=\sqrt{\frac{K_{t}}{K_{t-1}}}$; this would automatically map all the parameters from $\mathbb{L}^{n}_{K_{t-1}}$ to $\mathbb{L}^{n}_{K_{t}}$ and scale all distances between points linearly by that factor.
+As such, it preserves the relative distances between the hyperbolic parameters.
+However, the method itself presents issues during optimization and input scaling.
+Hyperbolic models suffer from instability and traditionally require higher precision computations to prevent invalid values.
+Mishne et al.
+[29] show that the mathematical instability is proportional to the distance between $\overline{\bm{0}}$ and the hyperbolic embeddings.
+They are also able to derive the maximum distance allowed before we begin to get undefined operations.
+By using the scaling factor $l$ we also scale $d_{\mathbb{L}}(\bm{x},\overline{\bm{0}})$, which could then push it outside the stable range.
+We theorize this is why works using this method need to adhere to strict training regimes, including burn-in periods without curvature learning, a separate optimizer for the curvature, and very low learning rates.
+Clipping these parameters is one solution, but larger $l$ values could make clipping too harsh, and it removes the nice property of preserved relative distances.
+Additionally, there is the issue of dealing with input projections, almost all hyperbolic learning approaches assume that the Euclidean or input data exists on the tangent plane of the origin and project it onto the manifold using the exponential map.
+This projection preserves the norm of the Euclidean vectors as $d_{\mathbb{L}}(\bm{x},\overline{\bm{0}})$.
+When the input is then projected onto the scaled manifold, the relative norms of the input and the hyperbolic class prototypes, for example, are now completely different, breaking their relationship.
+One would then have to find a way to scale the inputs while adhering to the maximum stability distance without clipping.
+This inconsistency is the same for mixed optimization settings where both Euclidean trivialized parametrization and direct hyperbolic parameter learning are used.
+An alternative mapping method is presented in Fu et al.
+[10], Guo et al.
+[15] and is based on projecting the parameters onto the tangent space at the origin $\overline{\bm{0}}_{t-1}$ of the old curvature using the logarithmic map and then projecting back after the curvature update.
+We show in Appendix D that this method preserves the distances and angles between the hyperbolic parameters and the origin.
+These properties are important as they are considered proxies for hierarchy level, and embedding similarity.
+We also show why the tangent space at the origin is the most mathematically suitable space for this mapping.
+This method also mitigates the instability caused by scaling, which removes the need for a more complicated training regime and maintains symmetric parameter handling by operating analogously to trivialized parameter learning.
+Given the above, our work relies on the tangent-based mapping method.
+We extend it to the optimization process by additionally defining the mapping of parameter gradients and momentums from $\mathcal{T}_{\bm{\theta}^{K_{t-1}}_{\mathbb{L}}}$ to $\mathcal{T}_{\bm{\theta}^{K_{t}}_{\mathbb{L}}}$.
+The entire mapping schema is shown in Algorithm˜1.
+We additionally perform empirical experiments to comparing the scaling mapping and the tangent mapping in Appendix E.
 It is important to emphasize that our proposed optimization scheme is compatible with existing curvature learning and meta-learning methods.
-Rather than being an alternative, it serves as a guideline for updating manifold parameters during curvature changes, aimed at maintaining learning stability throughout the process.
+Rather than being an alternative, it serves as an intermediate step for updating manifold parameters during curvature changes, aimed at maintaining learning stability throughout the process.
 3.3 Riemannian AdamW Optimizer
 Background
-The AdamW optimizer was first introduced by Loshchilov & Hutter (2019) and relies on an improved application of the L2-regularization factor in the base Adam optimizer.
+The AdamW optimizer was first introduced by Loshchilov and Hutter [26] and relies on an improved application of the L2-regularization factor in the base Adam optimizer.
 L2-regularization works on the principle that networks with smaller weights tend to have better generalization performance than equal networks with higher weight values.
 Adam applied the L2-regularization during the gradient update step by incorporating it in the loss.
-However, Loshchilov & Hutter (2019) argue that this is inconsistent since the regularization effect is reduced by the magnitude of the gradient norms.
+However, Loshchilov and Hutter [26] argue that this is inconsistent since the regularization effect is reduced by the magnitude of the gradient norms.
 Instead, they apply the weight decay directly during the parameter update step.
@@
 Given the above, we believe AdamW is significant for hyperbolic learning.
-Hyperbolic spaces, with their higher representational capacity, are prone to overfitting, especially under data scarcity during training (Gao et al., 2022).
+Hyperbolic spaces, with their higher representational capacity, are prone to overfitting, especially under data scarcity during training [12].
 Thus, AdamW’s enhanced L2-regularization could improve hyperbolic models.
@@
 The key difference between AdamW and Riemannian Adam lies in direct weight regularization, which is challenging in Lorentz space due to the lack of an intuitive subtraction operation.
-To address this, we re-frame parameter regularization in AdamW as a weighted centroid with the origin $\overline{\bm{0}}$
-$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda\overline{\bm{0}}$
+One option would be following the exponential mapping and retraction typically used for the optimization step.
+However we propose a simpler operation that reduces the need for expensive and less stable parallel transport operations.
+Specifically, we re-frame parameter regularization in AdamW as a weighted centroid with the origin
+$\bm{\theta_{t-1}}-\gamma\lambda\bm{\theta_{t-1}}=(1-\gamma\lambda)\bm{\theta_{t-1}}+\gamma\lambda\bm{O}$
 where $\gamma$ is the learning rate and $\lambda$ is the weight decay value.
-The centroid operation is well defined in the Lorentz space which now allows for a direct translation of AdamW.
-The regularization schema becomes:
-$\bm{\theta_{t}}=\begin{dcases}\mu_{\mathbb{L}}([\bm{\theta_{t-1}},\overline{\bm{0}}],\bm{\nu}=[1-\gamma\lambda,\gamma\lambda])&\quad\textrm{if }\bm{\theta}\in\mathbb{L}\\ \bm{\theta_{t-1}}-\gamma\bm{\theta_{t-1}}\lambda&\quad\textrm{otherwise}\end{dcases}$
-where $\mu_{\mathbb{L}}$ is the Lorentz centroid, and $\bm{\nu}$ are the weights. By removing the later gradient decay and introducing this operation, we adapt AdamW for use in the Lorentz space.
-Algorithm 1 Curvature Learning Aware Optimization
-  Given parameters $\bm{\theta}=[\bm{\theta}_{euclid},\bm{\theta}_{\mathbb{L}},\bm{\theta}_{K}]$ and gradients denoted $\mathcal{G}$
-  function Optimizer Step
-     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
-        Vanilla Riemannian Optimizer Step
-     end for
-     for $\bm{p}$ in $\bm{\theta}_{euclid}\cup\bm{\theta}_{K}$ do
-        Vanilla Euclidean Optimizer Step
-     end for
-     MoveParameters()
-  end function
-  function Move Parameters
-     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
-        $\bm{\mathcal{G}}_{temp}=\mathrm{PT}^{K_{t-1}}_{{\bm{p}}\rightarrow{\bm{\overline{\bm{0}}}_{t-1}}}\left({\bm{\mathcal{G}}}\right)$
-        $\bm{z}=\log_{\bm{\overline{\bm{0}}}_{t-1}}^{K_{t-1}}(\bm{p})$
-        $\bm{p}=\exp_{\overline{\bm{0}}_{t}}^{K_{t}}(\bm{z})$
-        $\bm{\mathcal{G}}=\mathrm{PT}^{K_{t}}_{{\bm{\overline{\bm{0}}}_{t}}\rightarrow{\bm{p}}}\left({\bm{\mathcal{G}}_{temp}}\right)$
-     end for
-  end function
+We can now directly translate this for hyperbolic parameters $\bm{\theta}_{\mathbb{L}}$ as $\mu^{\bm{\nu}}_{\mathbb{L}}([\bm{\theta_{t-1}},\overline{\bm{0}}])$ where $\mu^{\bm{\nu}}_{\mathbb{L}}$ is the weighted Lorentz centroid defined in Law et al.
+[22] and described in Appendix A.
+We define the centroid weights as $\bm{\nu}=[1-\gamma\lambda,\gamma\lambda]$.
+By removing the later gradient decay and introducing this operation as seen in Algorithm 2, we adapt AdamW for use in the Lorentz space.
+Algorithm 1 Tangent Based Manifold Mapping
+1:Given:
+2:Hyperbolic parameters $\bm{\theta}^{K}_{\mathbb{L}}$ on the $K$-curve manifold, Parameter gradients $\mathcal{G}\in\mathcal{T}_{\bm{\theta}^{K}_{\mathbb{L}}}$
+3:function Map Parameters
+4:  for each $\bm{p}\in\bm{\theta}_{\mathbb{L}}$ do
+5:   $\mathcal{G}_{\text{temp}}\leftarrow\mathcal{T}_{\bm{p}\rightarrow\overline{\bm{0}}_{t-1}}(\mathcal{G})$ $\triangleright$ Parallel transport gradient to previous origin
+6:   $\bm{z}\leftarrow\log^{K_{t-1}}_{\overline{\bm{0}}_{t-1}}(\bm{p})$ $\triangleright$ Project parameter onto tangent space
+7:   $\bm{p}\leftarrow\exp^{K_{t}}_{\overline{\bm{0}}_{t}}(\bm{z})$ $\triangleright$ Project back onto updated manifold
+8:   $\mathcal{G}\leftarrow\mathcal{T}_{\overline{\bm{0}}_{t}\rightarrow\bm{p}}(\mathcal{G}_{\text{temp}})$ $\triangleright$ Transport gradient back for next update
+9:  end for
+10:end function
+Algorithm 2 Riemannian Adam (RAdam) and Riemannian AdamW (RAdamW)
+1:Manifold $\mathcal{M}$, initial parameters $\bm{\theta}_{\mathbb{L}}\in\mathcal{M}$
+2:Learning rate $\alpha>0$, weight decay $\lambda\geq 0$, exponential decay rates $\beta_{1},\beta_{2}\in[0,1)$
+[…]
+4:Small constant $\epsilon>0$, max iterations $T$
+5:Optimized parameters $p_{t}\in\bm{\theta}_{\mathbb{L}}$
+6:Initialize moment vectors $m_{0}\leftarrow 0\in T_{p_{0}}\mathcal{M}$, $v_{0}\leftarrow 0\in T_{p_{0}}\mathcal{M}$
+7:Initialize timestep $t\leftarrow 0$
+8:for $t=1$ to $T$ do
+9:  $g_{t}\leftarrow\mathrm{grad}~f(p_{t-1})$ $\lambda\cdot p_{t-1}$
+10:  $\bm{\nu}\leftarrow[1-\gamma\lambda,\gamma\lambda]$
+11:  $p_{t-1}\leftarrow\mu^{\bm{\nu}}_{\mathbb{L}}([\bm{p_{t-1}},\overline{\bm{0}}])$
+12:  $g_{t}\leftarrow\text{egrad2rgrad}_{p_{t-1}}(g_{t})$ $\triangleright$ See Appendix A
+13:  $m_{t}\leftarrow\beta_{1}\cdot m_{t-1}+(1-\beta_{1})\cdot g_{t}$
+14:  $v_{t}\leftarrow\beta_{2}\cdot v_{t-1}+(1-\beta_{2})\cdot g_{t}\odot g_{t}$
+15:  $\hat{m}_{t}\leftarrow m_{t}/(1-\beta_{1}^{t})$
+16:  $\hat{v}_{t}\leftarrow v_{t}/(1-\beta_{2}^{t})$
+17:  $\eta_{t}\leftarrow\alpha\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
+18:  $p_{t}\leftarrow\text{Retract}_{p_{t-1}}(-\eta_{t})$ $\triangleright$ See Appendix A
+19:  $m_{t}\leftarrow\mathcal{T}_{p_{t-1}\to p_{t}}(m_{t})$
+20:end forreturn $p_{t}$
 3.4 Maximum Distance Rescaling
@@
 […]
-$D^{K}_{\overline{\bm{0}}_{max}}=arccosh(\frac{x_{t_{max}}}{\sqrt{K}})\cdot\sqrt{K}$ (2)
+$D^{K}_{\overline{\bm{0}}_{\text{max}}}=\text{acosh}\left(\frac{x_{t_{max}}}{\sqrt{K}}\right)\cdot\sqrt{K}$ (1)
 Under Float32 precision, and to account for values of $K<1$ we use a limit value of $x_{t_{max}}=2\cdot 10^{3}$.
-When projected onto the tangent space of the origin, and with $K=1$, this translates to $||\log^{1}_{\overline{\bm{0}}}(\bm{x})||=D^{1}_{\overline{\bm{0}}_{max}}=9.1$.
-This value changes considerably as the value of $K$ changes.
+When projected onto the tangent space of the origin, this translates to $\text{log}_{\overline{\bm{0}}}{\bm{x}}||_{\mathbb{E}}=D_{\overline{\bm{0}}_{max}}=9.1$.
 Vectors outside this radius lead to instability and performance degradation due to inaccurate approximation.
@@
 To constrain hyperbolic vectors within a specified maximum distance, either a normalization function or a parameter clipping method is required.
-Parameter clipping can be challenging to train, as it may lead to information loss and introduce non-smooth gradients.
+Parameter clipping can be challenging as it may lead to information loss and introduce non-smooth gradients.
 On the other hand, common normalization functions like tanh and the sigmoid function tend to saturate quickly, limiting their effectiveness as seen in the sigmoid implementation by Chen et al.
-(2022).
+[6].
 Flexible Scaling Function
 To address these issues, we introduce a modified scaling function, designed to provide finer control over both the maximum values and the slope of the curve.
-A visualization of this function is provided in Figure 2, and the formulation is presented below:
-$\bm{y}_{rescaled}=\frac{\bm{y}}{\lVert\bm{y}\rVert}\cdot m\cdot tanh(\lVert\bm{y}\rVert\cdot\frac{atanh(0.99)}{s\cdot m})$ (3)
-where $y\in\mathbb{R}^{d}$, $m$ is our desired maximum value, and $s$ controls the slope of the curve.
-We now have a maximum distance value to adhere to and a flexible distance normalizing function.
-To apply this to the hyperbolic vectors, we suggest performing the scaling on the tangent plane of the origin.
-However, this is an expensive operation to perform repeatedly, as such we derive in Appendix C the equivalent factorized form for the scaling of the space values:
-${\bm{x}_{s_{rescaled}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}}$ (4)
-where $D(\bm{x},\overline{\bm{0}})_{rescaled}^{k}$ is the scaled distances by plugging in $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Equation 3.
-The time component of the vector is recalculated based on the norm of its spatial components, providing a complete scaling operation for the Lorentz space.
-We apply this tanh scaling during transitions between manifolds, including from Euclidean to Lorentz spaces and between Lorentz spaces of different curvatures.
-It is also used after Lorentz Boosts and direct Lorentz concatenations (Qu & Zou, 2022), as well as following variance-based rescaling in the batch normalization layer, where adjustments can push points outside the valid radius, often resulting in invalid values.
-4 Domain Adaption
-In order to empirically prove the effectiveness of our proposed solutions, we apply them to metric learning, EEG classification, image classification, and image generation tasks.
-In the following, we tackle each of these tasks by offering a brief overview of the problem setting with an existing reference model, our hyperbolic adaptation for this problem, and the goal from this experimental setup.
+A visualization of this function is provided in Figure˜2, and the formulation is presented below:
+$\bm{y}_{\text{rescaled}}=\frac{\bm{y}}{\lVert\bm{y}\rVert}\cdot m\cdot\text{tanh}\left(\lVert\bm{y}\rVert\cdot\frac{\text{atanh}(0.99)}{s\cdot m}\right)$ (2)
+where $\bm{y}\in\mathbb{R}^{d}$, $m$ is our desired maximum value, and $s$ controls the slope of the curve.
+We now have a maximum distance value to adhere to and a flexible normalizing function.
+To apply this to the hyperbolic embeddings, we suggest performing the scaling on the tangent plane of the origin.
+However, this is an expensive operation to perform repeatedly, as such we derive in Appendix˜C the equivalent factorized form for the scaling of the space values:
+$\bm{x}_{s_{\text{rescaled}}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{\text{rescaled}}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{\text{rescaled}}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}$} (3)
+where $D(\bm{x},\overline{\bm{0}})_{\text{rescaled}}^{K}$ are the distances obtained by plugging $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Equation˜2.
+4 Experiments
+In order to empirically prove the effectiveness of our proposed solutions, we apply them to metric learning to test low precision learning, EEG classification for generalization ability in data hungry scenarios, and image classification and generation tasks for component efficiency.
+We also apply the curvature learning for most scenarios to study the new optimizer scheme and possible benefits.
+We include more detailed ablations and experiments on graph embeddings in the Appendix E.
 4.1 Hierarchical Metric Learning Problem
@@
 In their paper Kim et al.
-(2023) take on the problem of hierarchical clustering using an unsupervised hyperbolic loss regularizer they name HIER.
+[19] take on the problem of hierarchical clustering using an unsupervised hyperbolic loss regularizer they name HIER.
 This method relies on the use of hierarchical proxies as learnable ancestors of the embedded data points in hyperbolic space.
-Given a triplet of points ${xi,xj,xk}$ where $x_{i}$ and $x_{j}$ are determined to be related by a reciprocal nearest neighbor measure, and $x_{k}$ is an unrelated point, the HIER loss regularizer is then calculated as
-$\begin{split}\mathcal{L}_{\textrm{HIER}}(t)&=[D_{B}(x_{i},\rho_{ij})-D_{B}(x_{i},\rho_{ijk})+\delta]_{+}\\ &+[D_{B}(x_{j},\rho_{ij})-D_{B}(x_{j},\rho_{ijk})+\delta]_{+}\\ &+[D_{B}(x_{k},\rho_{ijk})-D_{B}(x_{k},\rho_{ij})+\delta]_{+},\end{split}$ (5)
-where $D_{B}$ denotes the hyperbolic distance on the Poincaré ball, and $\rho_{ij}$ is the most likely least common ancestor of points $x_{i}$ and $x_{j}$.
-This encourages a smaller hyperbolic distance between $x_{i}$, $x_{j}$, and $\rho_{ij}$, and a larger distance with $\rho_{ijk}$.
-The opposite signal is then applied in the case of $x_{k}$, the irrelevant data point.
-Kim et al.
-(2023) show substantial performance uplifts for the HIER loss when applied to a variety of network architectures.
-In the following experiment, we extend HIER to the Lorentz model (LHIER) and compare against the results provided by Kim et al. (2023).
+In the following experiment, we extend HIER to the Lorentz model (LHIER) and compare against the results provided by Kim et al.
+[19].
 Moving to Lorentz Space
-To adapt the HIER model to the hyperboloid, we first replace the Euclidean layer norm and linear layer with a Lorentzian norm and linear layer (Bdeir et al., 2024).
-We then modify the HIER loss by setting the hierarchical proxies as learnable hyperbolic parameters and replace the Poincaré distance with the Lorentzain distance.
-This allows us to optimize the hyperbolic parameters directly on the manifold using our new optimization schema.
-We also apply distance rescaling before layer norm and after the linear layer.
-We refer to this setup as LHIER.
+To adapt the HIER model to the hyperboloid, we first replace the Euclidean layer norm and linear layer with a Lorentzian norm and linear layer [3].
+We then modify the HIER loss by replacing the Poincaré distance with the Lorentzian distance and optimizing Lorentzian hierarchical proxy parameters directly on the manifold.
+We use our new optimization schema and apply distance rescaling before layer norm and after the linear layer.
 Experimental Goals
 Kim et al.
-(2023) employ the Euclidean AdamW optimizer with Euclidean parameterizations of hyperbolic proxies.
+[19] employ the Euclidean AdamW optimizer with Euclidean parameterizations of hyperbolic proxies.
 Their experiments use FP16 precision.
@@
 Any improvements likely stem from better curvature adaptation to the data and task.
-Experimental Setup
 We follow the experimental setup in Kim et al.
-(2023) and rely on four main datasets: CUB-200-2011 (CUB)(Welinder et al., 2010), Cars-196 (Cars)(Krause et al., 2013), Stanford Online Product (SOP)(Song et al., 2016), and In-shop Clothes Retrieval (InShop)(Liu et al., 2016).
-Performance is measured using Recall@k, the fraction of queries with at least one relevant sample in their k-nearest neighbors.
-All model backbones are pre-trained on ImageNet for fair comparisons.
+[19].
 Results
-Table 1: Performance of metric learning methods on the four datasets as provided by (Kim et al., 2023).
-All architecture backbones are pretrained and tested with the new LHIER loss.
-Superscripts denote their embedding dimensions and $\dagger$ indicates models using larger input images.
-As in (Kim et al., 2023), network architectures are abbreviated as, R–ResNet50 (Wang et al., 2017), B–Inception with BatchNorm (Ioffe & Szegedy, 2015), De–DeiT (Touvron et al., 2021), DN–DINO (Caron et al., 2021) and V–ViT (Dosovitskiy et al., 2020)
+Table 1: Performance of metric learning methods on the four datasets as provided by [19].
+$\dagger$ indicates models using larger input images.
+Network architectures are abbreviated as, R–ResNet50 [41].
 Methods Arch. CUB Cars SOP
@@
 CNN Backbone
-NSoftmax R128 56.5 69.6 81.6 88.7 75.2 88.7
-MIC R128 66.1 76.8 82.6 89.1 77.2 89.4
-XBM R128 - - - - 80.6 91.6
-XBM B512 65.8 75.9 82.0 88.7 79.5 90.8
-HTL B512 57.1 68.8 81.4 88.0 74.8 88.3
-MS B512 65.7 77.0 84.1 90.4 78.2 90.5
-SoftTriple B512 65.4 76.4 84.5 90.7 78.6 86.6
-PA B512 68.4 79.2 86.1 91.7 79.1 90.8
-NSoftmax R512 61.3 73.9 84.2 90.4 78.2 90.6
-ProxyNCA++ R512 69.0 79.8 86.5 92.5 80.7 92.0
-Hyp R512 65.5 76.2 81.9 88.8 79.9 91.5
-HIER R512 70.1 79.4 88.2 93.0 80.2 91.5
+NSoftmax [45] R512 61.3 73.9 84.2 90.4 78.2 90.6
+ProxyNCA++ [39] R512 69.0 79.8 86.5 92.5 80.7 92.0
+Hyp [9] R512 65.5 76.2 81.9 88.8 79.9 91.5
+HIER [19] R512 70.1 79.4 88.2 93.0 80.2 91.5
 LHIER R512 73.4 82.4 90.0 94.0 81.9 93.1
-Increase % 4.71 3.78 2.04 1.08 1.49 1.20
-As shown in Table 1, LHIER learns curvature without stability issues.
+Increase in % 4.71 3.78 2.04 1.08 1.49 1.20
+As shown in Table˜1, LHIER learns curvature without stability issues.
 Our approach improves model performance, with recall@1 gains ranging from 1.49% to 4.71%.
-Additional ablation experiments with different optimizers and configurations are detailed in Table 7.
+Additional ablation experiments with different optimizers and configurations are detailed in Table˜7.
 4.2 EEG Classification Problem
@@
 Each label is $y\in\{1,\dots,K\}$.
-Given $N$ labeled recordings $(X_{1},y_{1}),\dots,(X_{N},y_{N})$ from an unknown distribution $p$, the task is to train a model $\hat{y}$ that maps EEG signals $X$ to their correct class.
+Given $N$ labeled recordings $\left((X_{1},y_{1}),\dots,(X_{N},y_{N})\right)$ from an unknown distribution $p$, the task is to train a model $\hat{y}$ that maps EEG signals $X$ to their correct class.
 Reference Model
-The reference model MAtt (Pan et al., 2022) has already taken steps towards improving data preparation.
-In the feature extraction phase, MAtt uses 2 convolutions along the time and sensor dimensions respectively.
-The embeddings are transformed into Symmetric Positive Definite (SPD) data by mapping the feature embeddings from Euclidean space onto the SPD manifold.
-The resulting sequence is then fed into a Manifold Attention Model and the outputs are projected back to the Euclidean space and classified using a fully connected layer.
-Moving to Lorentz Space
-To adapt Matt to a Lorentz space setting, we project the embeddings from the SPD manifold to the Lorentz manifold after the input slicing and sample covariance matrix projection.
-We then pass them through an adapted Lorentz Attention layer that replaces the SPD layers, distance function, and centroid with the Lorentzian counterparts.
-Finally, we use the Lorentz MLR (Bdeir et al., 2024) as a final classification head.
-We refer to this enhanced model as HyperMAtt.
+In their work, Pan et al.
+[32] introduces Matt, a Riemannian model based on the SPD manifold.
+The model processes data through convolutional denoisers, extracts embedding covariances, and embeds them using a Riemannian attention mechanism before projecting the outputs back into Euclidean space for classification via a linear layer.
+Our proposed model, HyperMatt, extends this framework by projecting the resulting covariances onto the hyperboloid instead and processing them with a hyperbolic attention layer [6].
+The outputs are then classified directly on the manifold using the MLR [3].
 Experimental Goals
-While this process already represents EEG sequences in Riemannian space, we enhance it further with our Riemannian AdamW Optimizer.
 Hyperbolic spaces are prone to overfitting, especially in EEG classification, where models are trained per subject with limited sessions, leading to data scarcity.
-The proposed optimizer improves regularization, ensuring smoother and more balanced convergence.
-We avoid curvature learning, as it could worsen overfitting and reduce performance, but retain distance rescaling functions.
-Experimental Setup
-We compare our model HyperMAtt on three common EEG datasets for the domain of EEG classification.
-These datasets are MI (Motor Imagery) (Brunner et al., 2008), SSVEP (Steady-State Visual Evoked Potentials) (Nikolopoulos, 2021) and ERN (Error-Related Negativity) (Margaux et al., 2012) representing three different tasks, motor imagery, visual stimuli, and error recognition, respectively.
-Here, we follow the same preprocessing procedure and train/val/test splits as MAtt.
-For the datasets MI and SSVEP, we use Accuracy as an evaluation metric and AUC.
-For the application domain of EEG, we evaluate our method against current state-of-the-art models for EEG Classification including Inception (Burchert et al., 2024), MAtt (Pan et al., 2022), MBEEGSE (Altuwaijri et al., 2022), FBCNet (Mane et al., 2021), TCNet-Fusion (Musallam et al., 2021), EEG-TCNet (Ingolfsson et al., 2020), SCCNet (Wei et al., 2019), EEGNet (Lawhern et al., 2018), and ShallowConvNet (Schirrmeister et al., 2017).
+The proposed RAdamW optimizer theoretically improves regularization, ensuring smoother convergence.
+We apply our method to the datasets and the evaluation criteria in Pan et al.
+[32] to test this.
 Results
-In Table 2 we show the performance of HyperMAtt compared to the current state-of-the-art baselines, where the results were aggregated from (Pan et al., 2022; Burchert et al., 2024).
-The performance for HyperMAtt is measured across 3 runs and we report the mean and standard deviation.
-Here, our new optimizer archives state-of-the-art results for SSVEP and ERN, the two comparatively smaller datasets, which are more prone to overfitting.
-We also show the improved performance of our RAdamW optimizer vs the existing RAdam by training the model using both and comparing the results in Table 8.
+In Table˜2 we show the performance of HyperMAtt compared to the current state-of-the-art baselines, where the results were aggregated from [32, 5].
+The performance for HyperMAtt is measured across 10 runs and we report the mean and standard deviation.
+Here, our new optimizer achieves state-of-the-art results for SSVEP and ERN, the two comparatively smaller datasets, which are more prone to overfitting.
+We also show the improved performance of our RAdamW optimizer vs the existing RAdam by training the model using both and comparing the results in Table˜8.
 We also achieve these results in significantly less time/epoch compared to MAtt.
-Specifically for MI 0.077s/epoch vs 3.4s/epoch, for SSVEP 0.081s/epoch vs 4s/epochs, and for ERN 0.061s/epoch vs 1.9s/epoch.
+Specifically for MI 0.077s/epoch vs 3.4s/epoch, for SSVEP 0.081s/epoch vs 4s/epoch, and for ERN 0.061s/epoch vs 1.9s/epoch, resulting in an average speed-up by a factor of 42.
 Table 2: Performance comparison for the EEG datasets MI, SSVEP, and ERN. We report the average accuracy for MI and SSVEP and the AUC for ERN. The best result is highlighted in bold.
-Models MI SSVEP ERN ShallowConvNet 61.84$\pm$6.39 56.93$\pm$6.97 71.86$\pm$2.64 EEGNet 57.43$\pm$6.25 53.72$\pm$7.23 74.28$\pm$2.47 SCCNet 71.95$\pm$5.05 62.11$\pm$7.70 70.93$\pm$2.31 EEG-TCNet 67.09$\pm$4.66 55.45$\pm$7.66 77.05$\pm$2.46 TCNet-Fusion 56.52$\pm$3.07 45.00$\pm$6.45 70.46$\pm$2.94 FBCNet 71.45$\pm$4.45 53.09$\pm$5.67 60.47$\pm$3.06 MBEEGSE 64.58$\pm$6.07 56.45$\pm$7.27 75.46$\pm$2.34 Inception 62.85$\pm$3.21 62.71$\pm$2.95 73.55$\pm$5.08 MAtt 74.71$\pm$5.01 65.50$\pm$8.20 76.01$\pm$2.28 HyperMAtt (ours) 74.12$\pm$2.91 68.10$\pm$2.41 78.01$\pm$1.30 Increase in % -0.79 3.97 1.25
+Models MI SSVEP ERN
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Increase in % -0.78 4.01 2.59
 4.3 Standard Image Classification Problem
-Problem Setting and Reference Model
+Problem Setting and Model
 In their work, Bdeir et al.
-(2024) proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al.
-(2022).
-This approach ensured that the convolution outputs remained on the hyperboloid.
-However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions.
-As such, the authors proposed a two versions of their hyperbolic ResNet classifier.
-A fully hyperbolic model with all Lorentz ResNet blocks (HCNN) and a hybrid encoder model with alternating Euclidean and hyperbolic blocks (HECNN).
-Reference Optimization
+[3] propose a fully hyperbolic and hybrid encoder RessNet.
 To adapt our methods for both models, we introduce the distance rescaling function after every convolution in the encoder.
@@
 The end result is then equivalent to the convolution operation by Bdeir et al.
-(2024) while mitigating computation complexity.
-Specific implementation details can be found in Appendix B.
+[3] while mitigating computational complexity.
+Specific implementation details can be found in Appendix˜B.
 We denote our Hybrid and Fully hyperbolic models as HECNN+ and HCNN+ respectively.
 Experimental Goals
-In their paper, (Bdeir et al., 2024) notice lower performance in the fully hyperbolic models compared to the hybrid models and attribute this to instability in training.
+In their paper, [3] notice lower performance in the fully hyperbolic models compared to the hybrid models and attribute this to instability in training.
 Additionally, they cite crashes and model divergence when learning the curvature instead of setting it to constant.
 This would then be an ideal setting to test the optimization schema and rescaling function and their impact on performance and stability.
-We do not use the RAdamW optimizer in this setting as RSGD is already used and employs proper L2-regularization.
+Results
+In the ResNet-50 experiments in Table˜3, HECNN+ significantly outperforms both the Euclidean model and the base hybrid model across both datasets, demonstrating the positiv effect of curvature learning.
+We also see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime.
+We attribute this improvement to efficient closed-source CUDA convolution operations we can now leverage.
+One other benefit that we find from learning the curvature is quicker convergence, where the model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.
+ResNet-18 experiments show similar findings and can be found in Appendix E.
+Table 3: Performance and runtime Analysis for ResNet-50 models.
+We report classification accuracy (%) and the best performance is highlighted in bold (higher is better).
+All experiments were conducted on a single NVIDIA RTX 4090 GPU and an AMD EPYC 7543 CPU.
+CIFAR-100 Tiny-ImageNet VRAM tepoch
+$(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ For Cifar100 For Cifar100
+Euclid $78.52$ $66.23$ 4.5GB 30s
+HECNN 79.83{79.83} $66.30$ 15.6GB 300s
+HECNN+ $\bm{80.86}$ $\bm{67.18}$ 8.1GB 100s
+4.4 VAE Image Generation
 Experimental Setup
-We follow the experimental setup and hyperparameters of Bdeir et al.
-(2024) and rely on three main datasets: CIFAR10, CIFAR100, and Mini-Imagenet.
-For ResNet-18, we employ HECNN+, following the structure outlined in Bdeir et al.
-(2024) with alternating Euclidean and hyperbolic blocks, and integrate our efficient convolution operation.
-We decouple the encoder and decoder manifolds, allowing each to independently learn its own curvature for enhanced model flexibility.
-We also implement ResNet-50 models of the proposed hyperbolic hybrid architectures using similarly constructed bottleneck blocks.
-Results
-For the ResNet-18 experiments, Table 3 shows that the new models are able to remain stable while learning the curvature.
-Additionally, we see significant performance improvements between HCNN and HCNN+, where our proposed architecture now matches the performance of the hybrid encoders.
-We hypothesize that the improved scaling function helps mitigate the previous performance inconsistencies.
-However, the performance difference between the hybrid models is not significant, we hypothesize this is due to the alternating architecture which could limit the effect and instability of hyperbolic components.
-Finally, both models manage to maintain performance while reducing the memory footprint and runtime by approximately $25-50\%$ and $18-25\%$ respectively.
-One other benefit that we find from learning the curvature is quicker convergence where the model is able to reach convergence in 130 epochs vs the 200 epochs required by a static curve model.
-Table 3: Runtime Analysis for the ResNet-18 models.
-We report classification accuracy (%) and estimate the mean and standard deviation from five runs.
-The best performance is highlighted in bold (higher is better).
-CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ For Cifar100 $(\delta_{rel}=0.20)$ Euclidean $77.72_{\pm 0.15}$ 1.2GB - 12s $65.19_{\pm 0.12}$ Hybrid ($\mathbb{P}$) $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ Hybrid ($\mathbb{L}$) ${78.03_{\pm 0.21}}$ ${65.63_{\pm 0.10}}$ ResNet ($\mathbb{P}$) $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ HECNN ($\mathbb{L}$) ${78.76_{\pm 0.24}}$ 4.3GB - 100s ${65.96_{\pm 0.18}}$ HECNN+ ${78.80_{\pm 0.12}}$ 3GB - 80s ${65.98_{\pm 0.11}}$ HCNN ($\mathbb{L}$) ${78.07_{\pm 0.17}}$ 10GB - 175s ${65.71_{\pm 0.13}}$ HCNN+ $\bm{78.81_{\pm 0.19}}$ 5GB - 140s $\bm{66.12_{\pm 0.14}}$
-In the ResNet-50 experiments in Table 4, HECNN+ significantly outperforms both the Euclidean model and the base hybrid model across both datasets.
-We also see a $\sim 48\%$ reduction in memory usage and $\sim 66\%$ reduction in runtime.
-We attribute this improvement to efficient closed-source convolution operations we can now leverage.
-Table 4: Runtime Analysis for ResNet-50 models. We report classification accuracy (%) and the best performance is highlighted in bold (higher is better).
-CIFAR-100 VRAM/tepoch Tiny-ImageNet $(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ Euclid $78.52$ 4.5GB - 30s $66.23$ HECNN 79.8379.83{79.83}79.83 15.6GB - 300s $66.30$ HECNN+ $\bm{80.86}$ 8.1GB - 100s $\bm{67.18}$
-4.4 VAE Image Generation
-Problem Setting and Reference Model
-Previous research has highlighted the ability of hyperbolic neural networks to perform better when using lower dimensional embeddings due to the more expressive space (Nagano et al., 2019; Mathieu et al., 2019; Ovinnikov, 2019; Hsu et al., 2020).
-A natural extension to this would be implementing the models for VAEs which rely on smaller latent embedding dimensions to encode the inputs.
-Bdeir et al.
-(2024) perform this study on image generation and reconstruction for their fully hyperbolic models.
-Experimental Setup
-We reproduce the experimental setup from (Bdeir et al., 2024) and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers.
+We reproduce the experimental setup from [3] and re-implement the fully hyperbolic VAE using our new efficient convolution and transpose convolution layers.
 We also use curvature learning with our adjusted Riemannian SGD learning scheme.
 Results
-Our fully hyperbolic VAE implementation outperforms on both datasets, as shown in Table 5, while using 2.5x less memory and training 3x faster.
+Our fully hyperbolic VAE implementation outperforms on both datasets, as shown in Table˜4, while using 2.5x less memory and training 3x faster.
 This highlights the effectiveness of our curvature learning process and efficient model components.
@@
 These findings emphasize the importance of proper curvature adaptation in hyperbolic learning, paving the way for future research in optimizing hyperbolic embeddings across diverse fields.
-6 Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-We are committed to contributing positively to society and human well-being, while respecting the privacy of the subjects involved in our evaluation.
-Our experiments utilize three established EEG datasets: MI, SSVEP, and ERN, which involve data from human subjects.
-These datasets are fully anonymized and do not contain any personally identifiable information.
-Additionally, they are publicly available and widely used within the machine learning community.
+Further work should be done however to address the limitations on theoretical reasoning and efficiency presented in Appendix F.
+6 Acknowledgements
+Ahmad Bdeir was funded by the European Union’s Horizon 2020 research and innovation programme under the SustInAfrica grant agreement No 861924, as well as Horizon Europe Project BioMonitor4CAP grant agreement No 101081964.
 Appendix A Operations in hyperbolic geometry
@@
 A parallel transport operation $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)$ describes the mapping of a vector on the manifold $\bm{v}$ from the tangent space of $\bm{x}\in\mathbb{L}$ to the tangent space of $\bm{y}\in\mathbb{L}$.
-This operation is given as $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)=\bm{v}+\frac{\langle\bm{y},\bm{v}\rangle_{\mathbb{L}}}{K-\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}}(\bm{x}+\bm{y}).$
-Lorentzian Centroid (Law et al., 2019)
-Also denoted as $\bm{\mu}_{\mathbb{L}}$, is the weighted centroid between points on the manifold based on the Lorentzian square distance.
-Given the weights $\bm{\nu}$, $\bm{\mu}=\frac{\sum_{i=1}^{m}\nu_{i}\bm{x}_{i}}{\sqrt{1/K}\left|||\sum_{i=1}^{m}\nu_{i}\bm{x}_{i}||_{\mathcal{L}}\right|}.$
+This operation is given as $\mathrm{PT}^{K}_{{\bm{x}}\rightarrow{\bm{y}}}\left({\bm{v}}\right)=\bm{v}+\frac{\langle\bm{y},\bm{v}\rangle_{\mathcal{L}}}{K-\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}}(\bm{x}+\bm{y}).$
+Lorentzian Centroid [22]
+[…]
+Optimization Operations
+egrad2rgrad converts a Euclidean gradient (computed in the ambient space) to a Riemannian gradient by projecting it onto the tangent of the corresponding hyperbolic parameter.
+For the Lorentz manifold this is then:
+$\text{egrad2rgrad}(\mathbf{x},v)=v+\frac{\langle v,\mathbf{x}\rangle_{\mathcal{L}}}{K}\mathbf{x}.$
+The retraction maps the updated hyperbolic vector from the tangent space back onto the manifold. In the case of the hyperboloid this is the proposed exponential mapping equation.
 Lorentz Transformations
 In the Lorentz model, linear transformations preserving the structure of spacetime are termed Lorentz transformations.
-A matrix $\mathrm{\textbf{A}}^{(n+1)\times(n+1)}$ is defined as a Lorentz transformation if it provides a linear mapping from $\mathbb{R}^{n+1}$ to $\mathbb{R}^{n+1}$ that preserves the inner product, i.e., $\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{y}\rangle_{\mathbb{L}}=\langle\bm{x},\bm{y}\rangle_{\mathbb{L}}$ for all $\bm{x},\bm{y}\in\mathbb{R}^{n+1}$.
+A matrix $\mathrm{\textbf{A}}^{(n+1)\times(n+1)}$ is defined as a Lorentz transformation if it provides a linear mapping from $\mathbb{R}^{n+1}$ to $\mathbb{R}^{n+1}$ that preserves the inner product, i.e., $\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{y}\rangle_{\mathcal{L}}=\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}$ for all $\bm{x},\bm{y}\in\mathbb{R}^{n+1}$.
 The collection of these matrices forms an orthogonal group, denoted $\bm{O}(1,n)$, which is commonly referred to as the Lorentz group.
@@
 Specifically, in this context, Lorentz transformations satisfy the relation
-$\bm{O}^{+}(1,n)={\mathrm{\textbf{A}}\in\mathbb{R}^{(n+1)\times(n+1)}|\forall\bm{x}\in\mathbb{L}^{n}_{K}:\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{x}\rangle_{\mathbb{L}}=-\frac{1}{K},(\mathrm{\textbf{A}}\bm{x})_{0}>0)}.$ (6)
-Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ (Moretti, 2002).
+$\bm{O}^{+}(1,n)={\mathrm{\textbf{A}}\in\mathbb{R}^{(n+1)\times(n+1)}|\forall\bm{x}\in\mathbb{L}^{n}_{K}:\langle\mathrm{\textbf{A}}\bm{x},\mathrm{\textbf{A}}\bm{x}\rangle_{\mathcal{L}}=-\frac{1}{K},(\mathrm{\textbf{A}}\bm{x})_{0}>0)}.$ (4)
+Each Lorentz transformation can be decomposed via polar decomposition into a Lorentz rotation and a Lorentz boost, expressed as $\mathrm{\textbf{A}}=\mathrm{\textbf{R}}\mathrm{\textbf{B}}$ [30].
 The rotation matrix R is designed to rotate points around the time axis and is defined as
@@
 This structure shows that Lorentz rotations on the upper hyperboloid sheet belong to a special orthogonal subgroup, $\bm{SO}^{+}(1,n)$, which preserves orientation, with $\tilde{\mathrm{\textbf{R}}}\in\bm{SO}(n)$.
-In contrast, the Lorentz boost applies shifts along spatial axes given a velocity vector $\bm{v}\in\mathbb{R}^{n}$ with $||\bm{v}||<1$, without altering the time axis.
+In contrast, the Lorentz boost applies shifts along spatial axes given a velocity vector $\bm{v}\in\mathbb{R}^{n}$ with $\bm{v}||<1$, without altering the time axis.
 $\mathrm{\textbf{B}}=\left[\begin{array}[]{cc}\gamma&-\gamma\bm{v}^{T}\\ -\gamma\bm{v}&\mathrm{\textbf{I}}+\frac{\gamma^{2}}{1+\gamma}\bm{v}\bm{v}^{T}\end{array}\right],$ (6)
-with $\gamma=\frac{1}{\sqrt{1-||\bm{v}||^{2}}}$. However, this can also be any operation that scales the norms of the space values without changing the vector orientation.
+[…]
+(a) Tangent space at a random point on the manifold
+(b) Tangent space at the origin
+Figure 1: Tangent planes of a hyperboloid with curvature -1 relative to another hyperboloid with curvature -0.7.
+Tangential properties between manifolds are better respected at the origin where tangents remain parallel.
 Appendix B Convolution Trick
 To address the computational requirements issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al.
-(2021), which decomposes the transformation into a Lorentz boost and a Lorentz rotation.
+[8], which decomposes the transformation into a Lorentz boost and a Lorentz rotation.
 Using this definition, we replace the matrix multiplication employed by Bdeir et al.
-(2024) for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost.
+[3] for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost.
 Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
@@
 Finally, we formalize the new Lorentz Convolution as:
-$\bm{out}=\text{LorentzBoost(DistanceRescaling(RotationConvolution(}\bm{x}\textbf{)))}$)) (9)
+$\bm{out}=\text{LorentzBoost(DistanceRescaling(RotationConvolution(}\bm{x}\textbf{)))}$ (7)
 where TanhRescaling is the operation described in Eq.3 and RotationConvolution is a normal convolution parameterized through the procedure in Algorithm 3 where Transform is the norm-preserving transformation above.
-Algorithm 2 Lorentz Convolution Parametarization
-  $\bm{W}\in\mathbb{R}^{C_{in},C_{out},K_{width},K_{length}}$
-  function Adaptweight
-     if $K_{width}\cdot K_{length}\cdot C_{in}>C_{out}$ then
-        return $\bm{W}$
-     else
-        $\bm{W}=\bm{W}.reshape(K_{width}\cdot K_{length}\cdot C_{in},C_{out})$
-        $\bm{\hat{W}}_{core}=$ Transform ($\bm{W}$)
-        return $\bm{\hat{W}}.reshape(C_{in},C_{out},K_{width},K_{length})$
-     end if
-  end function
+Algorithm 3 Lorentz Convolution Parameterization
+[…]
+2:function AdaptWeight
+3:  if $K_{\text{width}}\cdot K_{\text{length}}\cdot C_{\text{in}}\leq C_{\text{out}}$ then
+4:   $\bm{W}\leftarrow\text{reshape}(\bm{W},K_{\text{width}}\cdot K_{\text{length}}\cdot C_{\text{in}},C_{\text{out}})$
+5:   $\bm{\hat{W}}_{\text{core}}\leftarrow\text{Transform}(\bm{W})$
+6:   $\bm{W}\leftarrow\text{reshape}(\bm{\hat{W}},C_{\text{in}},C_{\text{out}},K_{\text{width}},K_{\text{length}})$
+7:  end if
+8:  return $\bm{W}$
+9:end function
 Appendix C Scaling Lorentzian Vectors
@@
 $\bm{x}_{s_{rescaled}}=\bm{x}_{s}\times\frac{e^{\frac{D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})_{rescaled}^{K}}{\sqrt{K}}}}{e^{\frac{D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}-e^{\frac{-D(\bm{x},\overline{\bm{0}})^{K}}{\sqrt{K}}}}$ (11)
-Appendix D Ablations
-D.1 Image Classification
-We test the effect of individual model components in Table 6.
-Each subsequent model involves the default architecture presented in the experimental setup minus the mentioned component.
+Appendix D Proofs
+In the following, we show that the distances and angles between the origin and the hyperbolic points are preserved.
+This is relatively trivial but we add it for sake of clarity.
+For this discussion, we work in the Lorentz manifold and repeat the definitions here for easier referencing.
+$\mathcal{L}^{n}:=\{\mathbf{x}=[x_{t},\mathbf{x_{s}}]\in\mathbb{R}^{n+1}~|~\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=-K,~x_{t}>0\}$, $\mathbf{x},\mathbf{y}>_{\mathcal{L}}=-x_{t}y_{t}+\mathbf{x}_{s}^{T}\mathbf{y}_{s}$ and the origin is $O=[O_{0},\mathbf{O}_{s}]=[\sqrt{K},0,\dots,0]$.
+We also define two hyperboloids $H$ and $H^{\prime}$ with curvatures K-{K} and $K^{\prime}$} respectively.
+Let $x,y$ be two points on $H$ and $v,w\in T_{O}H$ be their logarithmic map onto the tangent space at the origin of $H$.
+Proof of distance preservation
+We can show that distances to the origin are preserved because $d_{\mathbb{L}}(\mathbf{x},\mathbf{O})$ collapses to $\lvert\lvert v\rvert\rvert_{euclid}$.
+This gives $d_{\mathbb{L}}(\mathbf{x},\mathbf{O})=\lvert\lvert v\rvert\rvert_{euclid}=d_{\mathbb{L}}(\mathbf{x^{\prime}},\mathbf{O^{\prime}})$.
+Proof of angle preservation
+We can similarly show that angles w.r.t to the origin are preserved. The angle $\theta$ between $x$ and $y$ at $O$ is actually computed via the Euclidean inner product in $T_{O}H$ using
+$\cos\theta\frac{<v,w>}{\lvert\lvert v\rvert\rvert\lvert\lvert w\rvert\rvert}$
+As such the angle w.r.t. origins also does not change when moving between manifolds of different curvature using this method since $v$ and $w$ don’t change.
+Appendix E Additional Experiments
+E.1 Image Classification
+Problem Setting and Reference Model
+In their work, Bdeir et al.
+[3] proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al.
+[6].
+This approach ensured that the convolution outputs remained on the hyperboloid.
+However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions.
+As such, the authors proposed a two versions of their hyperbolic ResNet classifier.
+A fully hyperbolic model with all Lorentz ResNet blocks (HCNN) and a hybrid encoder model with alternating Euclidean and hyperbolic blocks (HECNN).
+Table 5: Performance and runtime Analysis for the ResNet-18 models.
+We report classification accuracy (%) and estimate the mean and standard deviation from five runs.
+The best performance is highlighted in bold (higher is better).
+CIFAR-100 Tiny-ImageNet VRAM tepoch
+$(\delta_{rel}=0.23)$ $(\delta_{rel}=0.20)$ For Cifar100 For Cifar100
+Euclidean $77.72_{\pm 0.15}$ $65.19_{\pm 0.12}$ 1.2GB 12s
+Hybrid Poincaré [16] $77.19_{\pm 0.50}$ $64.93_{\pm 0.38}$ - -
+Hybrid Lorentz [3] $78.03_{\pm 0.21}$} $65.63_{\pm 0.10}$} - -
+Poincaré ResNet [40] $76.60_{\pm 0.32}$ $62.01_{\pm 0.56}$ - -
+HECNN [3]($\mathbb{L}$) $78.76_{\pm 0.24}$} $65.96_{\pm 0.18}$} 4.3GB 100s
+HECNN+ (ours) $78.80_{\pm 0.12}$} $65.98_{\pm 0.11}$} 3GB 80s
+HCNN [3] ($\mathbb{L}$) $78.07_{\pm 0.17}$} $65.71_{\pm 0.13}$} 10GB 175s
+HCNN+ (ours) $\bm{78.81_{\pm 0.19}}$ $\bm{66.12_{\pm 0.14}}$ 5GB 140s
+Resnet-18 Experiments
+Table˜5 shows that the new models are able to remain stable while learning the curvature.
+Additionally, we see significant performance improvements between HCNN and HCNN+, where our proposed architecture now matches the performance of the hybrid encoders.
+We hypothesize that the improved scaling function helps mitigate the previous performance inconsistencies.
+However, the performance difference between the hybrid models is not significant, we hypothesize this is due to the alternating architecture which could limit the effect and instability of hyperbolic components.
+Finally, both models manage to maintain performance while reducing the memory footprint and runtime by approximately $25-50\%$ and $18-25\%$ respectively.
+Table 6: Resnet-50 Ablations on Cifar-100.
+CIFAR-100 HCNN+ - Default $\bm{80.86}$ HCNN+ - fixed curve $79.6$ HCNN+ - no scaling $80.13$ HCNN+ - no optim scheme NaaNNaaN
+Ablation Experiments
+In table 6, we run ablation experiments to verify the effectiveness of our individual components.
+Specifically, the default case refers to the current model with the tanh rescaling function, learnable curvature, and our proposed optimization scheme.
+In the setting "fixed curve" we use a non-learnable curvature $K=1$, and keep the tanh rescaling.
+The use of the new optimization schema here has no effect since curvature is not learnable which means the staggered updates and parameter projections are not needed.
+In the setting "no scaling" we continue to learn the curvature with the new optimization schema but do not use the tanh rescaling.
+And in the setting "no optim scheme", we learn the curvature and use the scaling but do not use the new optimization schema.
 As we can see, the best results are achieved when all the architectural components are included.
 In the case of attempting to learn the curvature without the proposed optimizer schema, the model breaks completely down due to excessive numerical inaccuracies.
-Table 6: Ablation experiments for Resnet-50 models on Cifar-100.
-CIFAR-100 HCNN+ - Default $\bm{80.86}$ HCNN+ - fixed curve $79.6$ HCNN+ - no scaling $80.13$ HCNN+ - no optim scheme $NaaN$
-D.2 Metric Learning
-In Table 7 we present the LHIER model results under different optimizers and components.
+E.2 Metric Learning
+Problem Setting and Reference Model
+In their paper [19] rely on a hybrid hyperbolic architecture for modeling hierarchical relationships in image datasets.
+They include a new hyperbolic metric learning loss dubbed HIER which takes into consideration the intrinsic hierarchy in the data.
+Given a triplet of points xi,xj,xk{xi,xj,xk} where $x_{i}$ and $x_{j}$ are determined to be related by a reciprocal nearest neighbor measure, and $x_{k}$ is an unrelated point, the HIER loss is calculated as
+$\begin{split}\mathcal{L}_{\textrm{HIER}}(t)&=[D_{B}(x_{i},\rho_{ij})-D_{B}(x_{i},\rho_{ijk})+\delta]_{+}\\ &+[D_{B}(x_{j},\rho_{ij})-D_{B}(x_{j},\rho_{ijk})+\delta]_{+}\\ &+[D_{B}(x_{k},\rho_{ijk})-D_{B}(x_{k},\rho_{ij})+\delta]_{+},\end{split}$ (12)
+where $D_{B}$ denotes the hyperbolic distance on the Poincaré ball, and $\rho_{ij}$ is the most likely least common ancestor of points $x_{i}$ and $x_{j}$.
+This encourages a smaller hyperbolic distance between $x_{i}$, $x_{j}$, and $\rho_{ij}$, and a larger distance with $\rho_{ijk}$.
+The opposite signal is then applied in the case of $x_{k}$, the irrelevant data point.
+Kim et al.
+[19] show substantial performance uplifts for the HIER loss when applied to a variety of network architectures.
+We rely on four main datasets: CUB-200-2011 (CUB)[43], Cars-196 (Cars)[21], Stanford Online Product (SOP)[37], and In-shop Clothes Retrieval (InShop)[25].
+Performance is measured using Recall@k, the fraction of queries with at least one relevant sample in their k-nearest neighbors.
+All model backbones are pre-trained on ImageNet for fair comparisons.
+Ablation Experiments
+In table 7, we verify the effectiveness of our RAdamW optimizer, along with the additional components proposed.
+Specifically, "default" refers to the current LHIER model with learnable curvature and the new optimization scheme trained using RAdamW.
+The remaining settings are the same as default but differ in the mentioned component e.g.
+"RSGD" is trained with RSGD instead of RAdamW.
 Table 7: Ablation on components for Metric Learning
@@
 CNN Backbone
-LHIER RSGD R512 64.2 72.1 73.9 81.4 67.8 73.9
-LHIER RAdam R512 70.1 79.8 87.6 92.0 79.8 90.3
-LHIER CST Curvature R512 71.2 80.8 88.7 91.8 79.1 89.7
-LHIER No Learning Schema R512 65.8 72.0 - - - -
-LHIER R512 73.4 82.4 90.0 94.0 81.9 93.1
-D.3 EEG Classification
-We also train the model with the original Adam Optimizer and include the results in Table 8. This clearly shows the improvement provided by our RAdamW optimizer over standard RAdam.
+LHIER - RSGD R512 64.2 72.1 73.9 81.4 67.8 73.9
+LHIER - RAdam R512 70.1 79.8 87.6 92.0 79.8 90.3
+LHIER - Fixed Curvature R512 71.2 80.8 88.7 91.8 79.1 89.7
+LHIER - No Optim Scheme R512 65.8 72.0 - - - -
+LHIER - Default R512 73.4 82.4 90.0 94.0 81.9 93.1
+E.3 EEG Classification
+Ablation Experiments
+We also train the model with the original Adam Optimizer and include the results in Table˜8. This clearly shows the improvement provided by our RAdamW optimizer over standard RAdam.
 Table 8: Performance comparison for the EEG datasets MI, SSVEP, and ERN. We report the average accuracy for MI and SSVEP and the AUC for ERN. The best result is highlighted in bold.
-Models MI SSVEP ERN HyperMAtt + Adam (ours) 68,53$\pm$1.29 62.42$\pm$2.62 74.98$\pm$5.84 HyperMAtt + AdamW (ours) 74.12$\pm$2.91 68.10$\pm$2.41 78.01$\pm$1.30
+Models MI SSVEP ERN
+[…]
+[…]
+E.4 Graph Learning
+Problem Setting
+We follow the experimental setup described in Chen et al.
+[6] for the node classification problem in four popular graph embedding datasets.
+We choose the node classification task since the task decoder used by the fully hyperbolic graph convolution network (GCN) [6] relies on class prototypes learned directly on the hyperboloid.
+Additionally, the model is already Lorentz-based, this means we do not need to modify it, and we can test our components directly in a simple problem setting.
+Table 9 summarizes the training conditions for the ablations performed.
+Here, a checked AdamW refers to the use of the Euclidean AdamW for the trivialized model (hyperbolic parameters learned on the tangent space of the origin) and our proposed hyperbolic RAdamW for the others.
+We keep the same hyperparameters defined in Chen et al.
+[6] for all training settings.
+Additionally, for the scalemap-based method we scale the input data to match the new norms of the manifold before projecting.
+Results
+From table 9 we can see that that the model using Learnable curvature, our RAdamW and the tangent-based scaling achieves the highest performance.
+It should also be noted that the performance values for the the HyboNet paper are extracted directly from the paper [6], we were able to reproduce all results except the Disease dataset where we could only achieve a score of 91% using their setting.
+Table 9: Results on the node classification task using GCNs.
+Disease Airport PubMed Cora
+Trivialized AdamW Learnable K Scale Map
+HyboNet - - - - $96$ $90.9$ $78$ $80.2$
+H1 - - - $94.09$ $92.56$ $76.60$ $80.80$
+H2 - - - $92.89$ $93.25$ $77.23$ $81.5$
+H3 - - $94.82$ $93.6$ $78.3$ $82.1$
+H4 - $91.94$ $91.3$ $77.10$ $80.3$
+H5 - $94.88$ $93.55$ $78$ $81.8$
+Appendix F Limitations
+The proposed work still suffers greatly from the computational complexity introduced by hyperbolic operations.
+The effect of approximations for common operations such as the exponential map and the logarithmic map should be studied to reduce these issues.
+Additionally, alternatives for moving between manifolds should be researched and correlated with the task at hand and the loss being used.
+The different properties for the different alternatives could be better leveraged depending on the desired component outputs.